A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.03

Title word cross-reference

α [441]. L_{p} [68].

-Norm [68]. -Stable [441].

1 [48, 604]. 1750 [162]. 18th [649]. 1980s [424]. 1983 [468]. 1987 [12]. 1990s [113]. 1992 [343]. 1993 [218]. 1997 [649, 642, 730]. 19th [249]. 19th-Century [249]. 1st [12].

2 [598]. 2000 [145]. 2nd [278, 303, 408, 298].

3rd [297].

4 [264].

5th [12].

7th [164].

A. [294]. Abraham [263]. Abrupt [400]. Accelerated [18]. Accident [4]. According [583]. Accounting [745, 588, 489, 318, 423]. Accounts [224]. Acid [85, 312, 252, 617]. Acquired [316, 318, 688, 323, 320, 488, 708, 458]. Activities [245, 207]. Actuaries [469]. Adaptive [146]. Adelstein [263]. Adjusting [516]. Adjuvant [742]. Adult [716]. Adults [3]. Advanced [497, 604, 544, 376, 602, 604]. Advances [571]. Affairs [99]. Affect [718]. After [392]. Age [643, 114, 133, 681, 716, 584]. Age-related [584]. Age-Specific [133, 681].

Agency [172]. Agricultural [508]. Agriculture [187]. AIDS [331, 287, 431]. Aitken [593]. al [743, 748, 749, 744, 739, 740]. Albert [396]. Alcohol [253, 254, 4, 267]. Algebraic [412]. Algorithm [557]. Algorithms [531, 306]. Alldredge [415]. Allen [229]. Allocation [45, 225, 197]. Alsup [527]. Alternative [498]. ‘Alternatives [704, 603]. Altman [44]. America [186]. Among [153, 215]. Analysing [507, 449]. Analysis [278, 694, 526, 459, 424, 645, 489, 255, 486, 720, 741, 470, 458, 517, 220, 144, 15, 274, 644, 155, 344, 380, 389, 111, 145, 156, 269, 256, 290, 423, 125, 285, 83, 29, 543, 135, 536, 333, 576, 13, 406, 205, 210, 302, 414, 742, 69, 231, 521, 523, 413, 422, 573, 493, 531, 415, 440, 572, 516, 528, 52, 503, 353, 473, 416, 234, 238, 66, 529, 272, 241, 296, 571, 434, 102, 98, 506]. Analysts [48]. Analytic [448]. Analytical [375, 357]. Anderson [411, 528]. Annealing [101]. Annual [649]. Anonymized [115]. Ansell [11, 529]. Anthony [140]. Antitar [86]. Application [52, 483, 366, 488, 415]. Applications [586, 596, 162, 644, 574, 510, 376, 301, 413, 539, 531, 544, 409, 240, 439, 273, 400, 509, 266, 436, 531, 276]. Applied [25, 471, 437, 517, 235, 503, 473, 380, 296, 210, 302, 445, 446]. Appraisal [80, 43]. Approach [425, 59, 365, 484, 672, 707, 728, 575, 411, 431, 237, 24, 576, 330]. Approaches [304, 746, 363, 143]. Approximations [554, 651]. Arbitrary [738]. Area [484, 672, 731, 519]. Areas [103]. Armadale [346]. Armed [45]. Arnold [295]. Arrest [179]. Art [212]. Articles [501]. Artificial [221]. Aspects [101, 245, 207, 181, 81, 465, 208, 565]. Aspirations [215, 153]. Assessing [93]. Assessment [195, 85, 372, 388, 422]. Associated [285]. Associations [647]. Assume [186]. Assumptions [90]. Asymptotics [494]. Atkinson [121]. Atlas [597]. Atmospheric [543]. Attained [716].

Attitude [728]. Attitudes

[517, 164, 288, 124]. Attraction [305].

Attrition [520]. Audit [717].

Augmentation [750]. Austin [139, 393]. Author [670, 710]. Authority [103]. authors [710, 754]. Autocorrelated [224]. Automated [28]. Auxiliary [678]. Average [741]. Ayton [475].

B [63, 596, 334, 505, 470, 563, 530, 144, 273, 327, 601, 569, 42, 435, 64, 536, 24, 20, 576, 296, 295, 28, 567, 436, 236, 410, 493, 511, 382]. Babies [114]. Back [32, 53, 71, 108, 129, 148, 170, 190, 216, 246, 279, 313, 337, 358, 383, 419, 450, 476, 513, 545, 580, 606, 639, 673, 683, 698, 713, 724, 733, 757]. Backcalculation [319]. Background [77]. Bain [205]. Baines [396]. Bakeman [493]. Balakrishnan [183, 295, 534, 230]. Balanced [224]. Bandit [45]. Bangladesh [523]. Barker [563]. Barndorff [494, 465]. Barndorff-Nielsen [494, 465]. Barnett [466, 564, 495, 206]. Barnsley [428]. Bartholomew [595]. Base [741]. Base-Line [741]. Based [225, 483, 681, 550]. Basic [601]. Basis [482]. Basseville [400]. Batting [290]. Bayes [128]. Bayesian [473, 237, 184, 526, 92, 682, 746, 644, 722, 317, 708, 363, 12, 429, 264]. Bayesienne [244]. be [595, 480]. Becker [98]. Before [392, 162, 88]. Behavioral [229]. Behaviour [485, 331, 132, 441]. Behavioural [692]. Being’ [704]. Belief [595]. Belle [352]. Benjamin [327, 63, 202]. Bensaber [64]. Benzecri [231]. Beran [496]. Berg [401]. Berger [277, 264]. Bergman [505]. Bernardo [12, 429, 264]. Best [91, 501, 14, 748, 749]. Between [487, 321, 341, 647, 88]. Bias [588]. Bierens [497]. Billard [239]. Binary [520, 554, 651, 388]. Biologists [100]. Biology [568]. Biometry [301]. Biopharmaceutical [243, 276]. Biostatistical [271]. Biostatistics [352].

Birkes [498]. Bissell [430]. Bivariate [488]. Bleuse [64]. Bleuse-Trillon [64]. Blind [685]. Bohrer [371]. Book [329, 304, 122, 594, 624, 654, 418, 63, 278, 168, 402, 466, 544, 568, 50, 21, 23, 101, 356, 494, 235, 45, 596, 25, 121, 657, 27, 328, 209, 46, 443, 541, 626, 634, 667, 442, 625, 208, 433, 444, 575, 603, 185, 232, 334, 126, 51, 595, 43, 52, 141, 107, 438, 207, 308, 382, 448, 471, 534, 632, 267, 633, 655, 105, 411, 503, 437, 532, 469, 665, 303, 353, 505, 123, 147, 143, 307, 408, 409, 430, 470, 473, 239, 432, 563, 660, 298, 404, 530, 407, 144, 162, 635, 68, 31, 375]. Book [416, 15, 240, 351, 439, 566, 664, 167, 274, 565, 229, 242, 533, 373, 431, 527, 237, 434, 11, 128, 500, 230, 310, 265, 357, 70, 169, 658, 512, 380, 273, 327, 601, 102, 497, 638, 187, 145, 166, 574, 597, 271, 106, 627, 234, 188, 269, 331, 510, 165, 189, 243, 44, 48, 238, 417, 501, 535, 661, 12, 66, 99, 297, 372, 499, 630, 306, 564, 355, 502, 631, 668, 605, 186, 376, 529, 538, 103, 127, 125, 163, 593, 244, 441, 475, 569, 600, 467, 429, 104, 579, 16, 19, 29, 332, 543, 435, 629]. Book [636, 637, 669, 214, 663, 64, 536, 508, 22, 146, 212, 272, 381, 24, 333, 602, 309, 377, 20, 213, 277, 401, 161, 241, 542, 270, 268, 507, 576, 653, 301, 578, 628, 13, 296, 98, 400, 18, 540, 509, 49, 406, 183, 295, 412, 659, 28, 65, 205, 210, 299, 302, 330, 354, 403, 496, 233, 447, 414, 67, 567, 428, 571, 69, 405, 472, 379, 231, 604, 352, 266, 495, 17, 436, 47, 236, 164, 206, 211, 311, 184, 264, 335, 410, 445, 446, 498, 504, 599, 26, 30, 656, 413, 142, 275, 666, 573, 662, 465]. Book [539, 493, 378, 511, 531, 415, 449, 14, 440, 537, 570, 124, 276, 474, 100, 598, 572, 623, 305, 506, 374, 300, 468, 528, 577]. Bootstrap [300, 239, 435, 374, 577]. Boreham [399]. Boris [622]. Bose [8]. Bosworth [467]. Bounded [381]. Bourgeois [262]. Bourgeois-Pichat [262]. Bovine [549]. Bowden [99]. Bradford [139, 393, 392]. Branching [209]. Braun [468]. Breast [483, 742]. Breslow [13]. Brian [62]. Britain [706, 461, 361, 27, 597, 331, 211]. British [328, 343, 424, 730, 460, 482, 613, 116, 158, 164, 124]. Brockwell [65]. Brook [164, 124]. Brookmeyer [431]. Brown [402]. Brownlee [261]. BSc [41]. Buckley [207]. Building [719, 761, 107]. Bulmer [351]. Burden [619]. Burdick [265]. Business [596, 578]. Butler [14, 14]. Buuren [169]. Buydens [375]. Byar [201]. Byrne [530].

C [544, 356, 45, 121, 541, 603, 267, 503, 437, 532, 469, 307, 408, 404, 167, 373, 237, 357, 271, 269, 331, 510, 499, 355, 593, 244, 214, 20, 542, 18, 509, 183, 295, 414, 266, 445, 378, 531, 100, 598]. Calculus [376]. Calibration [402]. Campanelli [752, 753]. Campbell [100]. Campion [591]. Can [195, 82, 186, 456]. Cancer [745, 13, 742]. Capital [177]. Care [737]. Careers [37]. Carlo [599, 525]. Case [745, 382, 346, 480, 115]. Case-Study [346]. Cases [85, 312, 252, 323, 478, 616, 708, 705, 723]. Cash [587]. Castillo [122]. Casualties [727]. Categorical [735, 302]. Categories [738]. Catherine [348]. Causal [416, 43]. Causality [181]. CB [228]. CBE [350]. Celebration [25, 121]. Censored [266]. Census [732, 404, 110, 115, 361, 421, 158]. Censuses [351, 157]. Centenary [121]. Centile [682]. Central [172]. Century [2, 249]. Chain [707, 599]. Chakraborti [297]. Challenge [56]. Challenges [615]. Chance [58, 550]. Chandra [8]. Change [389, 457, 505]. Changes [282, 400, 218]. Changing [27, 727]. Chaos [58, 465]. Chaotic [441, 539]. Characterization [334]. Charles [592]. Charts [682]. Chatterjee [15]. Chaudhuri [16]. Chemistry [375, 357]. Chhikara [17]. Childhood [3, 705, 723]. Children [364]. Choe [416]. Choice [550, 242]. Choices [615]. Choquet [510]. Choquet-Deny [510]. Chow [544]. Christian [595, 370]. Chroniques [64]. Chui [531]. Churchill [427]. Circular [406]. City [179]. Clarifying [219]. Clarotti [18]. Class [485, 38, 2, 461, 214, 277]. Class-Party [461]. Classical [746]. Classification [648]. Classifications [676]. Classrooms [188]. Clayton [432, 752, 753]. Clear [615]. Cleveland [19, 433]. Clinical [278, 486, 155, 310, 372, 507]. Clinicians [323]. Clogg [499]. Clustering [586]. Clusters [94, 81]. Codman [717]. Coefficient [377]. Cofactors [689]. Cognitive [92]. Cohen [20, 183, 266, 305]. Cohort [13]. Coleman [328]. Collected [647, 468]. Collection [395, 218]. College [490]. Collett [403]. Comments [743, 748, 749, 744, 752, 739, 753, 740]. Commission [151]. Communicating [257]. Community [283]. Comparability [77]. Comparative [598, 23]. Comparison [741, 483, 721, 584]. Comparisons [677, 551, 301, 468]. Competitive [436]. Complementarity [232]. Completeness [323]. Completion [424]. Complex [526, 521]. Components [265, 163]. Compulsory [289]. Computational [101, 229]. Computer [57, 570, 241, 374]. Computer-Interactive [241]. Conceptions [736]. Conditional [585, 365]. Conference [614, 259]. Confidence [78, 75, 73, 35, 44, 34, 456, 265]. Conflict [146]. Conflux [146]. Confounders [345]. Conolly [596]. Consequences [222, 83]. Considerations [486]. Construction [489]. Consultant [136]. Consumer [648, 747]. Consumption [253, 254]. Contents [712, 756]. Context [620]. Contingency [525]. Continuous [389, 534]. Contraceptive [366]. Contribution [393]. Contributions [537]. Control [253, 375, 527, 22, 146, 28, 249, 537, 49]. Controlled [601]. Conventional [413].

Cooper [62]. Coordinate [128].

Coordinate-Free [128]. Correction [760]. Correlation [286, 330]. Correspondence [231]. Corrigendum [759, 245, 312, 672, 215, 761, 723]. Cottle [232]. Council [60, 137, 198, 291, 367, 462, 558, 649, 676]. Count [741, 344, 456, 504]. Counties [103]. Counting [78, 75, 353]. Countries [351, 695]. Counts [745, 459, 341]. Course [295]. Courses [173, 174]. Court [478, 64]. Courts [80]. Covariate [521]. Covariates [345, 287]. Coverage [46, 361]. Cows [550]. Cox [123, 500, 494, 500]. Creating [157]. Credit [648]. Cressie [233]. Cricket [290]. Crime [613, 89, 339, 113]. Criminal [615]. Criteria [601]. Critical [43]. Critique [197]. Crops [23]. Cross [310, 415]. Cross-Over [310, 415]. Crossed [750]. Crossings [536]. Crouchley [743, 744]. Cumulative [459]. Current [425]. Curriculum [93]. Curves [104]. Cycles [116]. Cyclic [570].

D

[568, 494, 119, 328, 541, 442, 595, 52, 438, 207, 448, 532, 469, 120, 353, 505, 123, 430, 432, 298, 416, 351, 566, 565, 229, 11, 512, 102, 145, 574, 597, 188, 269, 510, 189, 44, 501, 12, 297, 605, 103, 569, 600, 467, 309, 161, 241, 270, 268, 576, 18, 210, 299, 302, 403, 447, 352, 211, 335, 445, 498, 599, 413, 415, 14, 572, 305, 577]. Dagpunar [21]. Dale [404]. Daly [438, 532]. Daniel [9]. Darta [302]. Data [706, 647, 438, 455, 750, 741, 688, 220, 721, 520, 644, 155, 102, 366, 238, 620, 529, 686, 29, 543, 324, 607, 241, 507, 296, 735, 708, 210, 330, 403, 742, 521, 283, 449, 466, 433, 434, 125, 98, 406, 233, 504]. Dates [518]. David [347, 140, 349, 201]. Davies [743, 744]. Davis [65]. Dawid [264]. Dawson [307]. Day [13]. Daykin [469]. Dealing [709]. Death [341]. Decades [389]. Decayed [759, 729]. Decision [308, 458, 66, 83, 335, 474]. Decisions [312, 252]. Deconstructing [362]. Deficiency [316, 458, 318, 688, 323, 320, 488, 708]. Definition [583]. Definitions [582]. Degree [173, 174]. Degrees [178]. DeGroot [12, 40]. Dehnad [22]. Delay [318]. Delays [488]. Demands [93]. Deming [368]. Demographic [284, 255, 186]. Demography [166]. Densities [104]. Density [208, 309]. Dental [759, 729, 155]. Denti [145]. Deny [510]. Deoxyribonucleic [85, 312, 252, 617]. Dependence [645]. Depression [223]. Deprived [587]. Descriptive [571]. Design [278, 568, 694, 84, 486, 483, 696, 695, 423, 36, 22, 680, 13, 415, 537, 305, 563, 210, 440, 411, 236, 440]. Designs [570]. Detection [94, 557, 390, 400]. Determinants [707]. Developing [702, 351, 695]. Development [702, 608, 243, 467]. Developmental [518]. Developments [517]. Devolution [726]. DFC [228]. Diagnosis [316]. Diagnostic [489, 719, 761]. Dialetic [87]. Diamond [739, 740]. Dictionary [167]. Diet [643]. Different [483, 321]. Differential [718, 165]. Differentials [485]. Diffusions [541]. Diggle [434]. Direct [322]. Directive [283]. Disability [157]. Discrete [721, 600, 601]. Discrete-time [601]. Discriminant [503, 274, 333]. Discrimination [616]. Discussion [175, 481, 611, 704, 382]. Disease [646, 557, 135, 98, 689]. Diseases [94, 693]. Disorder [161]. Display [31]. Distortion [142]. Distribution [254, 307, 230, 17, 538]. Distribution-Free [538]. Distributions [442, 334, 534]. DNA [644, 691]. Do [194]. Doctor [112]. Dodge [498]. Does [718]. Doing [424]. Dorling [597]. Double [685]. Double-Blind [685]. Dr [62]. Draft [283]. Drawing [211]. Driver [340]. Drug [243, 134, 372]. Drugs [619]. Duffy [267]. Dugué [9]. Duncan [270]. Dunn [296].

Duration [721]. Dyke [23]. Dynamic [189, 19]. Dynamics [539, 321].

E. [398, 326]. Early [557, 214, 37].

Earnings [706, 679]. Eatwell [141]. Eckler [294]. Ecological [387]. Econometric [378]. Econometrics [143, 141, 497, 268]. Economic [304, 704, 245, 207, 11, 234, 103, 219, 676]. Economics [539]. Economies [142]. Economists [566]. Ecstasy [550]. Edgeworth [300]. Editor [589]. Editorial [701, 248, 642, 715, 700, 282, 478, 193, 549, 110, 132, 548, 218, 726, 219, 150, 453, 607, 641, 515, 675, 735, 685, 582, 676, 361, 172, 249, 339, 315, 283]. Editors [159, 258, 559, 621, 697, 650, 651, 199, 463, 226, 590]. edn [278, 303, 408, 298, 297]. Education [38, 2, 289]. Educational [738, 193]. Edward [62]. Edwards [368, 329]. Effect [221, 727, 750, 741, 522, 692, 90, 4, 339]. Effects [484, 672, 613, 680]. Efficacy [721]. Efron [435]. Eisenhart [427]. Elasticities [95]. Election [730, 641, 556, 343]. Elections [481]. Elementary [535, 447]. Elements [532, 184]. Elevated [365]. Elizabeth [10]. Ellickson [436]. Elwood [43]. Empirical [391, 107]. Employment [522, 583]. Enables [75]. Encephalopathy [549]. End [483]. Engelhardt [205]. Engineer [49]. Engineering [575, 277, 122]. Engineers [48, 572, 445]. England [284, 458, 688, 736, 322, 319, 519, 421, 321]. English [178]. Enthusiasm [709]. Enthusiasm-To-Respond [709]. Entrants [215, 153]. Entry [93, 38]. Environment [495, 564]. Environmental [341, 346, 703]. Environments [569]. Epidemic [316, 317]. Epidemics [324, 722]. Epidemiological [586, 267]. Epidemiology [759, 729, 555, 393, 431, 346, 705, 723, 432]. Epstein [533]. EQS [530, 530]. EQS/Windows [530]. Equation [530, 602]. Equations [510, 165]. Equilibrium [436].

Ermisch [234]. Erratum [672]. Error [678, 287, 516]. Errors [224]. Essentials [535, 268]. Estimate [322, 319]. Estimates [460, 519, 224, 583]. Estimating [95, 612, 340, 318, 4, 287, 681]. Estimation [619, 588, 643, 750, 746, 390, 731, 104, 309, 20, 678, 152, 183, 708, 388, 208, 68]. Ethics [618]. Ethnic [158, 179]. Europe [186]. European [283]. Evaluation [387, 593, 742]. Evans [415]. Evasion [142]. Event [524, 600]. Events [285, 518]. Everitt [296, 470]. Every [49]. Everywhere [428]. Evidence [706, 176, 718, 92, 85, 80, 616, 82, 610, 593, 617, 67, 550]. Evidence-Based [550]. Evolving [116]. Exact [525, 196]. Examination [490, 487]. Examining [647, 484, 672]. Exchangeability [220]. Executives [277]. Exercise [256, 422, 50]. Expansion [300]. Expectations [166]. Expenditure [610]. Experiment [737]. Experimental [568, 563, 210, 511, 440, 537, 566]. Experiments [23, 411, 36, 415, 440, 102, 508]. Expert [618, 18]. Expertise [479]. Explaratory [269]. Exploration [751]. Exploratory [238]. Exploring [239]. Exponential [612, 681]. Exposure [340]. Extension [340]. Extreme [122].

F [329, 443, 185, 438, 532, 303, 144, 167, 565, 527, 500, 265, 234, 269, 12, 564, 355, 429, 301, 28, 428, 495, 264]. Faced [482]. Factor [380]. Fallacy [286]. Family [389, 610]. Fang [405]. Farms [11]. Fatal [4]. Federal [615]. Feedback [401]. Feige [142]. Felons [113]. Female [178]. Feminization [679]. Fewer [113]. Field [23, 508]. Fielding [739, 740]. Fields [512, 69]. Fienberg [121, 748]. Filipino [364]. Filled [759, 729]. Finance [282]. Financial [378]. Fingerprinting [85]. Finite [236]. First [106, 295]. FIS [41]. Fisher [352, 136, 406, 196]. Fitting [488]. Fixed [285]. Fleming [437, 353, 598]. Flenberg [749]. Florence [347]. Florens [184]. Flows [165, 678]. Flury [24]. Folks [17]. Follow [588]. Follow-up [588]. Following [218]. Food [332]. Foods [95]. Football [221]. Forecasting [304, 730, 409, 473, 578, 317]. Forecasts [585]. Forensic [614, 386, 644, 620, 593]. Formal [127]. Formed [482]. Forms [273]. Formula [225]. Forwards [675]. Foundations [541]. Four [155]. Fowler [652]. Fox [242, 214]. Fractal [305, 305]. Fractals [512, 428]. Fractional [681, 719, 761]. Frame [613]. Framework [391, 607]. Frank [426]. Frederick [652]. Free [128, 538]. Freeman [565]. French [66]. Frequencies [483]. Frequency [504]. Friedman [566]. Front [1, 33, 55, 72, 109, 131, 149, 171, 192, 217, 247, 281, 314, 338, 360, 384, 420, 452, 477, 514, 547, 581, 640, 674, 684, 699, 714, 725, 734, 758]. FSS [41]. Fuels [482]. Full [522]. Full-Time [522]. Functional [510]. Functioning [484, 672]. Functions [306, 488, 518]. Fundamentals [271, 269]. Future [173, 186]. Fuzzy [574].

G [304, 544, 23, 541, 105, 437, 505, 143, 409, 298, 375, 439, 274, 380, 102, 187, 574, 331, 44, 372, 593, 475, 508, 212, 333, 602, 161, 270, 507, 296, 98, 299, 604, 352, 445, 326, 14, 598]. Gail [431]. Galambos [471]. Gambling [533]. Games [99]. Gani [25, 471]. Gardner [44, 293]. Garthwaite [567]. Gastwirth [67]. Gauss [128]. Gaussian [17]. Geisser [407]. Gender [425, 679, 589]. General [481, 649, 211, 556, 343, 70, 472]. Generalized [518, 126]. Generated [570]. Generation [21]. Genstat [379]. Geoffrey [397]. Geometric [128]. Geometry [605, 573]. George [228]. Geospatial [607]. Gibbons [297]. Gilbert [372]. Giles [11]. Gilks [599]. Giri [373]. Gittins [45]. Given [93]. Glasserman [600]. Gnedenko [622].

Godfrey [143]. Gonin [68]. Government [282, 389, 219, 676, 283]. Graduates [37]. Graduation [255]. Graphical [395, 707]. Graphics [19]. Graybill [265]. Great [211, 706, 361]. Green [501]. Grenander [472]. Grimmett [298, 161, 299]. Groenendaal [356]. Gross [678]. Group [158]. Groups [321]. Growth [116]. Guesswork [615]. Guide [163, 598, 404]. Guidelines [44]. Gujarati [268]. Guttorp [209].

H [304, 418, 27, 603, 68, 118, 439, 167, 431, 512, 497, 187, 166, 574, 165, 12, 125, 543, 272, 24, 295, 330, 567, 275, 537, 468]. Hadi [15]. Haemophilia [690]. Haigh [611]. Hald [162]. Hall [46, 300]. Halmos [128]. Hamaker [370]. Hammersley [161]. Hand [438, 532]. Handbook [438, 470, 277, 231, 230]. Handled [619]. Hannagan [26]. Hannan [398]. Hansen [439, 118]. Härdle [235]. Harmonization [77, 582]. Harold [491]. Harrington [353]. Harrison [473]. Harry [591]. Harvey [408]. Hazards [424]. Health [364, 647, 484, 672, 423, 737, 150, 3, 352, 197, 194, 67, 214]. Heath [568]. Hedayat [236]. Helen [350]. Herman [203]. Hernández [601]. Hernández-Lerna [601]. Herschel [152]. Heterogeneity [692]. Higher [38, 536]. Highest [716]. Hill [139, 392, 393]. Hills [432]. Hinkelmann [440]. Historical [394]. Histories [524]. History [162, 166]. Hjorth [374]. Hoaglin [269]. Hoffmann [502]. Hoffmann-Jorgensen [502]. Holgate [325]. Home [221]. Honour [471, 161]. Honours [173]. Hope [482]. Hoppe [301]. Housing [180, 361]. Howson [237]. Hubele [28]. Huberty [503]. Hughes [569]. Hugo [370]. Human [316, 687, 694, 588, 690, 692, 322, 696, 695, 691, 99, 317, 708, 319, 689, 321]. Hundred [649]. Hurwitz [439]. Hypothesis [391].

Identifiability [334]. Identification [386, 644, 146]. Ignatov [50]. Ignorable [709]. II [502, 29]. III [696]. Illness [485, 114, 267]. Illustrated [128]. ILOG [493]. Image [573]. Immigration [85]. Immune [316, 458, 318, 688, 323, 320, 488, 708]. Immunization [523]. Immunodeficiency [316, 687, 694, 588, 690, 692, 322, 696, 695, 691, 317, 708, 319, 689, 321]. Impact [680]. Implicating [617]. Imprisonment [339]. Improved [554, 651, 480]. Improvement [342]. Improving [112]. Inaccuracies [745]. Incidence [588, 727, 318, 59, 133]. Income [95, 679, 91, 180]. Inconsistency [487]. Incorrect [312, 252]. Independent [700, 540]. Index [670, 710, 756, 754, 759, 729, 105, 336]. India [95, 176, 249]. Indicators [304, 704, 702, 490, 5, 703]. Indices [45]. Individual [176]. Individuals [690]. Induced [340, 319]. Industrial [103, 116]. Industries [177, 332, 275]. Industry [596]. Inequalities [2, 3]. Inequality [679]. Infant [341, 707]. Infected [690]. Infection [694, 322, 708, 689]. Infectious [557, 98]. Inference [386, 209, 455, 387, 407, 722, 183, 236, 185, 303, 271, 297, 567, 494]. Inflated [759, 729]. Influence [727, 692, 215, 153]. Information [316, 54, 130, 191, 280, 359, 451, 546, 671, 711, 755, 741, 458, 193, 708, 142, 31]. Innovative [75]. Inspector [91]. Institutional [677, 551]. Insurance [51]. Intake [643]. integrating [703]. Integrity [35, 34]. Intensity [488]. Intensive [374]. Interactions [284]. Interactive [241]. International [245, 207, 77, 188, 12, 127, 124, 598, 614, 583, 598]. Interplay [688]. Interpretations [449]. Intervals [265, 44, 681, 584]. Intervention [695]. Interviewer [680]. Interviewers [751]. Introduction [79, 568, 46, 218, 435, 542, 440, 407, 189, 572, 373, 417, 272].

Introductory [444]. Inverse [17]. Investigating [586]. Investigation [341, 391]. Involved [478]. Isham [123]. ISI [121]. Issues [694, 702, 478, 77, 551, 696, 482, 602, 495]. Item [728]. Ivanov [69].

J [278, 168, 402, 50, 21, 101, 356, 45, 25, 328, 442, 208, 232, 126, 595, 43, 141, 107, 438, 207, 471, 267, 411, 503, 437, 532, 307, 473, 144, 274, 229, 434, 11, 601, 497, 234, 188, 269, 331, 44, 48, 12, 99, 297, 355, 502, 529, 538, 163, 429, 16, 332, 543, 435, 214, 146, 381, 333, 213, 401, 161, 28, 65, 205, 210, 302, 354, 496, 414, 67, 571, 398, 231, 17, 47, 184, 264, 335, 504, 326, 599, 26, 465, 511, 415, 570, 598, 506, 374, 577, 390]. J-Shaped [390]. J. [369]. J.-M [184]. J.-P [231, 184]. J.-S [232]. Jackknife [577]. Jackson [163]. Jaffe [47]. Jambu [238]. James [491]. Jan [560]. Janacek [409]. Janicki [441]. January [218]. Jean [350, 262]. Jennett [492]. Jensen [465]. Jewish [460]. Jobson [210, 302]. John [161, 570, 468]. Johnson [534, 102, 535]. Jolliffe [567]. Jones [532, 20, 330, 567]. Joreskog [380]. Jorgensen [502, 410]. Joseph [492, 41]. Joshi [27]. Jowell [164, 124]. Judge [91]. June [12, 649, 642]. Jurek [442]. Jurisdiction [76]. Jurors [82]. Just [535].

K. [261]. K.-T [405]. K.-Y [434]. Kaleidoscope [596]. Kallaa [562]. Kalton [270]. Karr [185, 303]. Kasprzyk [270]. Kateman [375]. KCB [399]. Keats [28]. Kedem [536]. Keep [195]. Kelly [443]. Kempthorne [440]. Kendall [354, 465]. Kenneth [592]. Khuri [376]. Kiers [125]. Kingdom [207, 307, 187, 467, 332, 245, 134]. Kingman [355]. Kipling [249]. Kleijnen [356]. Klein [275]. Know [49]. Knox [331]. Kotz [534]. Krzanowski [571]. Kunita [165].

L [541, 334, 534, 411, 505, 143, 409, 239, 375, 434, 125, 467, 42, 401, 205, 67, 352, 17, 164, 142, 465, 511, 531, 14, 124, 474]. Laarhoven [101]. Labour [701, 394, 583]. Lagrange [143]. Lahiri [304]. Lajos [471]. Lampeter [382]. L’Analyse [244]. Lancaster [166]. Large [750, 423]. Lasserre [601]. Latent [517, 728]. Law [259, 67]. Layouts [506]. Leading [304]. League [551]. Least [603]. Lee [305]. Legal [81, 87, 83]. Leitch [572]. Lemeshow [240]. Lending [111]. Lenin [6]. Leonenko [69]. LePage [239]. Lerna [601]. Lessons [717]. Letter [159, 258, 559, 621, 697, 650, 651, 199, 226, 590, 589]. Letters [463]. Leukaemia [705, 723]. Level [727, 692, 742]. Levy [240]. Lewis [48, 466, 542]. Liang [434]. Lieshout [573]. Lievesley [211]. Life [195, 166, 20, 18, 205, 703, 51]. Lifestyles [288]. Light [727]. Likelihood [329]. Limit [442, 540]. Limitations [551, 250]. Limits [239]. Lindley [565, 12, 18, 565]. Lindsey [504]. Line [741, 537]. Linear [544, 232, 126, 15, 518, 410, 493, 603]. Linkage [57]. Lists [746]. Litigation [482, 705, 723]. Loans [111]. Local [255, 282, 103]. Log [493]. Log-Linear [493]. Logic [533]. Logistic [520, 230]. Logothetis [537]. London [555, 113]. Lone [234]. Long [284, 95, 645, 496, 716]. Long-Memory [496]. Long-Range [645]. Long-Term [284, 716]. Longford [377, 739, 740]. Longitudinal [718, 741, 505, 434, 742, 330]. Look [106]. Looking [116, 675]. Lorenzen [411]. Lottery [608, 609, 610]. Low [180]. Luck [479]. Lunn [438, 532, 241]. Lutkepohl [272]. Lutz [186]. Lynn [211].

M [50, 101, 43, 141, 437, 532, 469, 473, 432, 530, 416, 351, 439, 565, 242, 431, 527, 128, 500, 70, 273, 574, 44, 238, 417, 12, 529, 127, 429, 579, 19, 161, 270, 507, 576, 301, 400, 205, 414, 428, 184, 264, 573, 415, 474, 598].

MacArthur [331]. Machin [576]. Macintosh [305]. Macroeconomic [154]. Macroeconomics [381]. Mad [550]. Madow [439, 439]. Magnusson [505]. Maintaining [73]. Malaysia [707]. Malnutrition [176]. Managers [277, 542]. Managing [277]. Mandel [506]. Manie [263]. Manly [144]. Mann [444]. Manor [214]. Manton [574]. Manual [50, 379]. Manufacturing [527, 28]. Mapping [745]. Marazzi [306]. Marcoulides [602]. Mardia [212]. Maritz [538]. Mark [332]. Markers [689]. Market [701, 218]. Markov [541, 70, 601, 335, 599, 474]. Marriott [167]. Marsh [348, 404]. Martin [293]. Martingales [541]. Marubini [507]. Mason [442]. Mastering [47, 26]. Matched [721]. Matching [57]. Mathematical [418, 596, 178, 472, 321]. Mathematics [173, 604, 51]. Matrices [489]. Matter [32, 53, 71, 1, 33, 55, 108, 129, 148, 72, 109, 131, 170, 190, 216, 149, 171, 192, 246, 279, 313, 217, 247, 281, 337, 358, 383, 314, 338, 360, 419, 450, 476, 384, 420, 452, 513, 545, 580, 477, 514, 547, 606, 639, 673, 581, 640, 674, 683, 698, 713, 684, 699, 714, 724, 733, 757, 725, 734, 758]. Matthai [273]. Matusita [29]. Maurice [354, 592]. McConway [438, 532]. McCullagh [126]. McGill [19]. McLachlan [274, 333]. McNeese [275]. McNeil [241]. Mean [518, 487]. Means [395, 731]. Measure [157]. Measured [287]. Measurement [454, 642, 425, 485, 193, 552, 177, 678, 224, 589, 516, 402]. Measures [690]. Measuring [382, 693, 132, 617]. Medhi [213]. Medical [392, 403]. Medicine [43]. Medicines [548]. Medio [539]. Meeting [649, 704, 12, 34, 649]. Meier [357]. Memory [496]. Men [487]. Messy [102]. Meta [720]. Meta-Analysis [720]. Metcalfe [575]. Method [340, 322, 681, 22]. Methodological [732]. Methodology [688, 75, 48, 352, 288]. Methods [586, 387, 505, 430, 240, 439, 566, 357, 648, 529, 125, 135, 13, 405, 498, 275, 511, 374, 584, 538, 183, 65, 205, 302, 206, 446]. Michael [41]. Might [480]. Milestones [518]. Milgate [141]. Miller [168]. Milliken [102]. Mills [378]. Mining [455]. Minorities [179]. Mirazchüski [50]. Miss [350]. Missing [759, 729, 620, 521]. Misspecification [143]. Mixed [518]. Mixing [321]. Mixture [612]. Mobility [718, 553, 222]. Model [759, 729, 107, 455, 391, 518, 374, 418]. Modeling [530, 147]. Modelling [364, 645, 760, 59, 365, 646, 484, 672, 721, 461, 488, 154, 707, 602, 735, 708, 747, 378, 403, 504]. Models [544, 526, 603, 334, 750, 57, 612, 720, 741, 432, 416, 738, 520, 554, 366, 499, 651, 324, 213, 728, 412, 205, 388, 719, 761, 573, 449, 321, 126, 408, 510, 417, 377, 20, 414, 410]. Modular [604]. Moir [307]. Money [68]. Monitoring [316, 382, 194]. Monotone [600]. Monte [599, 525]. Montefusco [531]. Montgomery [445]. Monthly [583]. Months [114]. Moore [304, 611]. Morbidity [485]. Mormon [156]. Moroney [41]. Morris [118, 40]. Mort [103]. Mortality [718, 690, 285, 707, 716, 166, 327]. Mosteller [269]. Mouchart [184]. Move [327]. Multi [45]. Multi-Armed [45]. Multidimensional [500]. Multigroup [147]. Multilevel [484, 672, 746, 554, 188, 686, 651, 751, 742, 388, 523, 516]. Multilocus [312, 252]. Multipaper [413]. Multiple [746, 272, 468, 301]. Multiplicity [486]. Multiplier [143]. Multistate [366]. Multivariable [128, 719, 761]. Multivariate [155, 238, 605, 731, 296, 210, 302, 571, 24, 309]. Mysterious [81].

N [308, 534, 439, 373, 230, 510, 377, 13, 98, 406, 183, 295, 28, 233, 69, 573, 537]. Nadaraya [104]. Nagaraja [295]. Name [411]. Nation [150]. National [704, 718, 700, 608, 607, 224, 288, 609, 610, 151, 197]. Natural [380]. Nature [223]. Need [225, 194]. Needs [74]. Nelder [126]. Nellis [437, 598]. Networks [693, 412, 465]. Neumann [175]. Newman [141]. Nicholson [127]. Nielsen [494, 465]. Nigerian [177]. Night [730]. Nightingale [347, 717]. Nikiforov [400]. Nineties [56]. No [411]. No-Name [411]. Non [709, 102, 6, 36, 751, 728, 51]. Non-Ignorable [709]. Non-Life [51]. Non-Orthogonal [36]. Non-Replicated [102]. Non-Response [709, 751, 728]. Non-Soviet [6]. Nonlinear [68]. Nonparametric [235, 208, 446, 297, 104]. Norm [68]. Normal [681]. Norms [92]. North [186]. Note [320, 315]. Notices [336]. Nuclear [482]. Number [732, 105, 405]. Number-Theoretic [405]. Nuss [145]. Nutrient [643]. Nutrients [95].

O [494, 601, 166, 214, 264, 465, 440]. Obituary [139, 371, 347, 561, 119, 397, 350, 140, 202, 120, 349, 348, 10, 9, 8, 426, 336, 399, 118, 201, 369, 560, 368, 370, 492, 97, 40, 42, 491, 562, 62, 396, 427, 228, 398, 591, 592, 622, 652, 41, 464, 204, 203]. Observational [507]. Observed [722]. Occupation [647]. Occupational [425, 589]. Occurrence [391]. Ocean [543]. Odeh [242]. Off [537]. Off-Line [537]. Offending [449]. Office [718, 157, 172, 583]. Official [78, 394, 73, 35, 180, 34, 76]. Oldham [119]. Oliveira [369]. O’Muircheartaigh [752, 753]. One [649, 732]. One-Number [732]. Operations [48]. Operator [442]. Operator-Limit [442]. Opinion [481, 645, 480, 18, 556]. Opportunity [215, 153]. Optimal [308, 169]. Optimality [601]. Optimization [443]. Orav [48]. Ord [354]. Order [536, 295, 183]. Ordered [738]. Ordinal [499, 742]. Organization [76]. Orthogonal [36]. Ostrowski [438]. Other [143]. Outbreaks [557]. Outcomes [193]. Outliers [686, 466]. Overall [157].

P [418, 402, 101, 356, 209, 46, 443, 444, 126, 141, 207, 353, 240, 351, 565, 237, 434, 357, 48, 475, 600, 270, 578, 412, 65, 567, 231, 211, 184, 264, 446, 30, 537, 300]. P. [97]. Paine [228]. Palgrave [141]. Panel [520, 706, 270]. Pang [232]. Papers [175, 611, 471, 743, 748, 753, 740, 749, 744, 752, 739]. Parameter [20]. Parental [38]. Parenthood [234]. Parmar [576]. Partially [722]. Participation [178, 610, 289]. Partner [693]. Partnerships [487]. Party [461, 197]. Paste [86]. Paternity [85, 312, 252]. Pattern [274, 333, 728, 472, 716]. Patterns [328, 135, 154]. Pattison [412]. Payer [91]. Payne [379]. Peace [243, 276]. Pearce [207]. Peery [201]. Penrice [397]. Pentikainen [469]. Percival [413]. Peremptory [615]. Performance [221, 677, 551, 490, 256, 742, 5]. Personal [393]. Perspective [394, 133, 356, 188]. Pesonen [469]. Peter [443, 562, 119]. Peters [187]. Petersen [508]. Petrov [540]. Phase [696]. PhD [424, 37]. Phenomenon [59]. Philip [325]. Philippines [176]. Phillips [529]. Physical [161, 742, 413, 511]. Physics [511]. Pichat [262]. Picturing [343]. Pitch [221]. Pitman [326]. Pittenger [229]. Pius [562]. Plaintiff [88]. Planned [308]. Planning [578]. plc [482]. Plus [470]. Point [59, 365, 646, 512, 285, 185, 303, 123]. Points [483, 738]. Poisson [759, 729, 355]. Polderman [146]. Pole [473]. Policy [253, 67]. Political [645, 760, 222, 479]. Poll [645]. Polling [479]. Polls [481, 480, 641, 556]. Polynomials [681, 719, 761]. Popularity [645, 760]. Population [585, 745, 27, 328, 741, 690, 746, 692, 460, 157, 236, 361, 519, 211, 421, 14, 63]. Populations [144, 240, 250, 194, 147]. Port [509]. Position [642]. Post [289].

Post-Compulsory [289]. Potential [692, 90]. Poverty [679]. Power [214]. Practical [575, 43, 189, 24, 576, 469, 527, 529]. Practice [309, 599]. Prakasa [334]. Pratique [64]. Pre [641]. Pre-Election [641]. Precensus [284]. Precision [680]. Predictability [251]. Predicted [727]. Prediction [251]. Predictions [225, 458]. Predictive [407]. Predictors [719, 761]. Presence [678]. Present [77]. Presentation [395]. presented [649]. Prespecified [59, 285]. Prevalence [322, 133, 319]. Prévision [64]. Primary [737]. Primer [566, 277]. Principal [163]. Principle [143]. Principles [77, 206, 21, 437]. Prior [708]. Probabilistic [82, 465]. Probabilities [252, 617, 312]. Probability [50, 442, 334, 471, 162, 373, 417, 83, 104, 480, 540, 509, 445, 25, 475, 447, 443, 448, 298, 502, 299]. Probes [312, 252]. Problem [232, 105]. Problems [425, 299, 505]. Procedure [87]. Procedures [301, 388, 308]. Proceedings [12]. Process [59, 365, 527, 28, 275]. Processes [541, 185, 303, 353, 601, 106, 299, 328, 209, 46, 123, 298, 70, 355, 441, 496, 335, 474]. Processor [401]. Producing [583]. Production [116]. Products [747]. Professor [120, 348, 42, 293, 203]. Profiling [644]. Prognostic [719, 761]. Programmes [721]. Programming [189, 418]. Progress [738]. Progressing [287]. Progression [128, 689]. Project [482]. Projection [255]. Projections [688, 320, 145]. Proof [619, 87]. Prophylactic [692]. Proportions [731]. Prospective [215, 153]. Protection [283, 703]. Provincial [179]. Provision [737]. Provost [273]. Public [73, 111, 35, 34, 67, 5, 556]. Puccio [531]. Pul [414]. Pupils [188]. Pure [604]. Puri [97]. Puterman [474]. Pyzdek [277, 49].

Quadratic [273]. Qualitative [125]. Quality [248, 342, 195, 527, 277, 49, 249, 537, 703, 382, 563, 375, 22, 277, 537]. Quality-of-Life [195]. Quantify [114]. Quantitative [677, 31, 431, 125]. Quantities [619]. Question [158]. Questions [362]. Queueing [213]. Queues [401].

R [494, 603, 232, 52, 107, 207, 353, 123, 147, 239, 298, 68, 31, 416, 565, 242, 533, 431, 527, 11, 265, 357, 380, 106, 510, 535, 99, 467, 579, 332, 435, 508, 602, 277, 161, 241, 65, 299, 330, 379, 17, 164, 599, 275, 493, 415, 449, 570, 124, 100, 572]. R. [136]. Raised [59]. Raj [8]. Random [21, 750, 741, 298, 512, 273, 569, 540, 518, 299, 69, 106, 569, 377]. Random-Effect [750, 741]. Randomized [363]. Range [645, 681]. Rao [603, 510, 334]. Rare [94, 59]. Rasbash [752, 753]. Rates [424, 588, 88, 116, 179]. Ratio [286]. Rational [595]. Rationality [381]. Ratkowsky [415]. Raudenbush [188]. Readings [66]. Realistically [526]. Reasoning [237, 67]. Reay [482]. Recalled [518]. Receiving [742]. Recognition [274, 333]. Reconciling [518]. Record [57]. Records [57, 115, 304]. Refereeing [685]. Reference [682, 379, 681, 584]. Refinements [677]. Regional [320]. Regions [103]. Registration [745]. Regression [402, 341, 52, 646, 520, 104, 210, 197, 521, 168, 235, 15, 498]. Regular [472]. Regularity [254]. Regulation [548, 134]. Regulations [84]. Related [156, 495, 584]. Relation [646]. Relations [127]. Relationship [341, 88, 390, 461]. Relationships [43, 643]. Relative [680]. Release [379]. Reliability [529, 20, 205, 414, 18, 572]. Religious [553]. Rendell [11]. Reoffending [612]. Replicated [102]. Report [60, 137, 198, 259, 291, 367, 462, 558, 649, 164, 124]. Reporting [318, 91, 323, 488, 487].

Research [687, 505, 351, 482, 256, 467, 13, 676, 422, 310, 403]. Residence [484, 672]. Resource [197, 197]. Resources [225]. Respond [709]. Responding [74]. Response [709, 390, 751, 728]. Responses [554, 651, 388]. Responsibilities [76]. Restart [522]. Results [490, 256]. Retherford [416]. Return [522]. Revealed [643]. Reverse [366]. Review [329, 304, 122, 624, 654, 418, 63, 278, 168, 402, 466, 544, 568, 50, 21, 23, 101, 356, 494, 235, 45, 596, 25, 121, 657, 27, 328, 209, 46, 443, 541, 626, 634, 667, 442, 625, 208, 433, 444, 575, 603, 185, 232, 334, 126, 51, 595, 43, 52, 141, 107, 438, 207, 308, 382, 448, 471, 534, 632, 267, 633, 655, 105, 411, 503, 437, 532, 469, 665, 303, 353, 505, 123, 147, 143, 307, 408, 409, 430, 470, 473, 239, 432, 563, 660, 298, 404, 530, 407, 144, 162, 635, 68, 31, 375, 416]. Review [15, 240, 351, 439, 566, 664, 167, 274, 565, 229, 242, 533, 373, 431, 527, 237, 434, 11, 128, 500, 230, 310, 265, 357, 70, 169, 658, 512, 380, 648, 273, 327, 601, 102, 497, 638, 187, 145, 166, 574, 597, 271, 106, 627, 234, 188, 269, 331, 510, 165, 189, 243, 44, 48, 238, 417, 501, 535, 661, 12, 66, 99, 297, 372, 499, 630, 306, 564, 355, 502, 631, 668, 605, 186, 376, 529, 538, 103, 127, 125, 163, 593, 244, 441, 475, 569, 600, 467, 429, 104, 579, 16, 19, 29, 332, 543, 435, 629]. Review [636, 637, 669, 135, 214, 663, 64, 536, 508, 22, 146, 212, 272, 381, 24, 333, 602, 309, 377, 20, 213, 277, 401, 161, 241, 542, 270, 268, 507, 576, 653, 301, 578, 628, 13, 296, 98, 400, 18, 540, 509, 49, 406, 183, 295, 412, 659, 28, 65, 205, 210, 299, 302, 330, 354, 403, 496, 233, 447, 414, 67, 567, 428, 571, 69, 405, 472, 379, 676, 231, 604, 352, 266, 495, 17, 436, 197, 47, 236, 164, 206, 211, 311, 184, 264, 335, 410, 445, 446, 498, 504, 599, 26, 30, 656, 413, 142, 275, 666, 573]. Review [662, 465, 539, 493, 378, 511, 531, 415, 449, 14, 440, 537, 570, 124, 276, 474, 100, 598, 572, 623, 305, 506, 374, 300, 468, 528, 577]. Reviews [594, 624, 654, 245, 207, 307, 187, 467, 332]. Revisited [286]. Reyment [380]. Ribaudo [743, 744]. Richard [204, 464]. Richardson [599]. Riedwyl [24]. Right [111]. Risk [612, 469, 340, 365, 646, 390, 549, 4, 321]. Risk-Response [390]. Risks [287, 30]. Road [727, 4]. Robert [371, 244]. Robinson [493]. Robust [306, 22]. Rogers [541]. Role [687, 136, 176, 751]. Rolin [184]. Romaniuk [748, 749]. Ronald [652]. Ross [294, 417]. Routinely [647]. Routines [306]. Royal [649, 34]. Rudinger [505]. Rules [92]. Run [95]. Runger [445]. Rural [95, 523].

S [544, 596, 121, 433, 444, 232, 334, 534, 120, 470, 407, 15, 240, 566, 434, 310, 169, 273, 327, 145, 188, 417, 66, 297, 372, 499, 306, 538, 97, 19, 212, 277, 296, 509, 17, 236, 164, 311, 599, 465, 511, 124, 374, 470]. S-Plus [470]. S.-C [544]. Safe [549]. Safety [372]. Salmond [349]. Salt [328]. Same [549]. Sample [439, 680, 439, 242, 206]. Samples [115, 266, 211]. Sampling [715, 144, 423, 480, 613, 16, 236, 240, 311]. Sargent [381]. SAS [47]. Scales [728]. Scaling [169, 500]. Scenarios [585]. Schmitz [308]. Schoen [147]. School [93, 490]. Schooling [188]. Schools [215, 153, 188]. Schumacker [602]. Science [92, 212, 37, 30, 511, 596]. Sciences [229, 380, 579, 543, 352]. Scientific [237]. Scientist [344]. Scientists [593]. Scope [219]. Scores [738]. Scoring [114, 648]. Scotland [317, 708, 289]. Scott [10, 309]. Scottish [178, 215, 153]. Screening [483]. Scripture [156]. Search [158]. Seasonal [154]. Secondary [389]. Sector [5]. Segregation [425, 589]. Selecting [345, 14]. Selection [168, 301, 250, 374]. Selectivity [720, 256]. Sellafield [705, 723]. Semiparametric [643]. Senn [310].

Sensitivity [489, 458, 15]. Sequences [540]. Sequential [278, 308, 276]. Serial [330]. Series [645, 341, 408, 409, 473, 501, 536, 154, 272, 65, 169, 354, 378, 528]. Service [700, 283, 197]. Services [76]. Session [60, 137, 198, 291, 367, 462, 558, 649]. Sets [438, 574]. Severe [319]. Severity [157]. Sexual [693, 132, 288, 487, 331]. Sexually [693]. Shanbhag [510]. Shao [577]. Shaped [390]. Shapes [512]. Sharing [401]. Sharpe [70]. Shearer [578]. Shecter [277]. Shihadeh [499]. Short [95, 587]. Should [49]. Shrinkage [731]. Sidney [202]. Silvey [120]. Simons [331]. Simple [195]. Simulated [101]. Simulation [409, 144, 356, 48, 441]. Singh [270]. Single [218]. Sinha [236]. Sir [139, 393, 399, 591, 204]. Sirkin [579]. Sixty [649]. Sixty-third [649]. Size [718, 746, 242]. Skills [56]. Small [438, 11, 731, 519]. Smith [565, 189, 12, 429, 264]. Smithers [604]. Soccer [221]. Social [718, 38, 489, 222, 517, 344, 597, 103, 579, 457, 3, 412, 676, 37, 164, 124, 351]. Society [7, 39, 61, 96, 117, 138, 160, 182, 200, 227, 260, 292, 550, 649, 34]. Socioeconomic [718]. Software [229, 414]. Sojourn [401]. Soliman [327]. Solutions [299]. Solving [418, 89]. Some [586, 392, 181, 81]. Son [382]. Source [63]. Sources [245, 207, 307, 646, 187, 467, 332, 708, 598]. South [95, 736]. Soviet [6]. Space [330]. Span [20]. Spatial [586, 736, 135, 233, 573]. SPC [430]. Special [124]. Specific [133, 681]. Spectral [413]. Speigelhalter [599]. Spongiform [549]. Spooner [542]. Sprent [446, 30]. Spurious [286]. Square [525]. Squares [603]. Stability [505]. Stable [522, 441]. Stage [144]. Stage-Structured [144]. Standard [286]. Standards [84]. Stanford [511]. State [330]. Statistical [649, 466, 356, 136, 185, 245, 455, 207, 255, 57, 181, 303, 700, 702, 307, 470, 557, 86, 75, 81, 80, 167, 274, 533, 90, 616, 527, 193, 551, 696, 362, 648, 111, 187, 145, 256, 44, 297, 133, 290, 538, 34, 467, 332, 135, 212, 151, 333, 76, 582, 172, 446, 717, 422, 465, 283, 276, 374, 584, 528, 209, 430, 432, 416, 357, 574, 99, 499, 29, 543, 13, 406, 28, 205, 414, 67, 567, 69, 275, 511]. Statistician [6]. Statisticians [478, 48, 548, 134]. Statistics [701, 112, 687, 7, 39, 61, 78, 96, 117, 138, 160, 182, 200, 259, 248, 227, 260, 292, 63, 392, 568, 704, 385, 121, 443, 718, 342, 394, 223, 162, 618, 77, 73, 128, 257, 344, 609, 552, 35, 166, 691, 243, 180, 501, 12, 502, 548, 34, 218, 726, 219, 24, 134, 542, 183, 604, 264, 250, 445, 89, 194, 598, 174, 614, 444, 437, 532, 373, 535, 306, 605, 376, 19, 295, 405, 184, 26, 573, 575, 229, 44, 564, 593, 579, 496, 233, 495, 100]. Statistique [244]. Status [172]. Stefan [561]. Stirzaker [298, 299, 447]. Stochastic [334, 510, 165, 441, 722, 165, 213, 573]. Stone [232, 128, 464, 204]. Story [86]. Stoyan [512, 512]. Stoyanov [50]. Strange [332]. Strategies [16]. Strategy [732]. Straub [51]. Strength [617]. Strong [617]. Stroock [448]. Structural [602, 530]. Structure [517, 600]. Structured [735, 144]. Structures [472]. Students [424]. Studies [745, 188, 507, 13, 382, 471]. Studio [382]. Study [718, 690, 721, 166, 346, 179, 11]. Stuvel [105]. Stylometric [156]. Subject [336, 598]. Subjective [83, 475]. Subset [168]. Success [88]. Sudden [341]. Suicide [555]. Summertime [727]. Sums [459]. Sunder [566]. Surface [221]. Surgical [717]. Surrogate [483]. Survey [706, 439, 16, 157, 680, 751, 206, 521, 288, 610, 613]. Surveys [715, 351, 389, 423, 157, 270]. Survival [353, 366, 507, 403, 576]. Sustainability [703]. Sustainable [702]. Swift [409]. Symmetric [728]. Syndrome [316, 341, 318, 688, 323, 320, 488, 708, 458]. Syski [106]. System [43, 114, 305, 47].

Systems [161, 600].

T [466, 411, 469, 353, 563, 271, 605, 381, 377, 277, 49, 567, 405, 26, 413, 378, 528]. Tables [551, 525, 14]. Taguchi [22, 537]. Takacs [471]. Taking [30]. Tanushev [50]. Tarling [449]. Tax [91, 142]. Tax-Payer [91]. Teachers [501]. Teaching [93, 501]. Team [221]. Techniques [602, 413]. Teenage [736]. Teeth [759, 729]. Telephone [715]. Temperature [341]. Term [284, 549, 716]. Terme [64]. Terms [167]. Test [196]. Testing [489, 18, 205]. Tests [143, 525]. Text [459]. Texts [156]. Their [185, 303, 162, 551, 83]. Theorems [540]. Theoretic [405]. Theoretical [101, 509]. Theories [127]. Theory [122, 544, 46, 93, 448, 469, 439, 533, 70, 552, 273, 87, 16, 29, 309, 400, 540, 65, 205, 472, 266, 436, 410, 539, 531, 50, 442, 429, 213]. Therapeutic [694]. Therapy [742]. Thiebaux [543]. Third [614, 351, 12, 649]. Thompson [350, 107, 311]. Three [125]. Three-Way [125]. Tiago [369]. Tibshirani [435]. Time [706, 341, 473, 522, 721, 169, 154, 272, 378, 528, 408, 409, 536, 65, 354, 601]. Times [401]. Tinbergen [560]. Toby [228]. Today [186]. Tolley [574]. Tool [527]. Tooth [86]. Tooth-Paste [86]. Topics [497]. Torestad [505]. Tort [67]. Toutenberg [603]. TQM [430]. Trade [218]. Transformation [681, 719, 761]. Transitions [366, 518]. Transmission [693, 321]. Transmitted [693]. Treatment [180]. Trends [392, 328, 186, 103, 457]. Trial [483, 88, 618, 696]. Trials [694, 486, 615, 155, 310, 695, 507, 363, 278, 372]. Tribute [443, 565, 212]. Trillon [64]. Truncated [266]. Tu [577]. Tufte [31]. Tukey [468, 269]. Turkman [564, 495]. Tutor [229]. Tutorial [229]. Tweedie [592]. Twentieth [2]. Two [389, 125, 506]. Two-Way [125, 506]. Type [692, 510].

U [472, 374]. UK

[63, 207, 103, 454, 642, 553, 732, 245, 715, 690, 679, 344, 256, 726, 154, 608, 641, 556, 224]. Uncertain [225, 595]. Uncertainty [585, 455, 565, 30]. Unchecked [90]. Under-Reporting [91]. Underground [142]. Understand [82]. Understanding [112, 493]. Unemployed [721]. Unemployment [454, 642, 344, 583]. Unified [16]. United [207, 307, 187, 467, 332, 245, 134]. Univariate [534]. Universities [178, 256]. University [215, 153]. untitled [657, 626, 634, 667, 625, 632, 633, 655, 665, 660, 635, 664, 658, 638, 627, 661, 630, 631, 668, 629, 636, 637, 669, 663, 653, 628, 659, 656, 666, 662, 623]. Updating [682, 519]. Uptake [523]. Urbach [237]. Urban [194]. Usage [177]. Use [316, 738, 366, 747, 319, 5]. User [74, 404, 163]. Using [459, 647, 612, 709, 312, 252, 470, 746, 490, 574, 391, 678, 719, 761, 322].

V [466, 23, 575, 411, 123, 565, 12, 564, 212, 400, 18, 540, 69, 495, 206]. Vaccine [696]. Vaccines [692]. Vajda [561, 596]. Valencia [12]. Validation [374]. Validity [35, 34]. Valsecchi [507]. Value [122, 37]. Values [521]. Vardeman [511]. Variable [709, 728]. Variables [273, 678, 499, 540]. Variance [265, 518, 269]. Variate [21]. Variation [176, 736]. versus [482]. Very [11]. Vicinity [59]. View [393, 502, 6, 448]. Viewpoint [267]. Views [283]. VIII [468]. Virus [316, 687, 694, 588, 690, 692, 322, 696, 695, 691, 317, 708, 319, 689, 321]. Viry [145]. Visual [31]. Visualization [524, 309]. Visualizing [433]. Vol [541, 245, 207, 307, 439, 102, 187, 48, 569, 467, 332, 13, 210, 302, 67, 511, 531, 440, 598, 468]. Vols [502]. Volume [54, 130, 191, 280, 359, 451, 546, 670, 671, 711, 755, 161, 710, 712, 756, 754, 121]. Vos [16].

W [329, 356, 235, 433, 232, 448, 439, 188, 269, 48, 186, 16, 19, 146, 309, 277, 571, 379, 599, 275, 465, 468, 528]. W. [368]. Wagner [391]. Wakefield [748, 749]. Walden [413]. Wales [458, 688, 322, 319, 519, 421, 321]. Walks [569, 569]. Wang [544, 405]. Warwick [351]. Water [495]. Watson [527, 212]. Wavelet [531]. Wavelets [531]. Way [125, 506, 125]. Weighted [459]. Welch [42]. Well [704]. Well-Being’ [704]. Welsh [161]. Weron [441]. West [473, 736]. Westover [604]. White [335]. Whitehead [278]. Whitten [20]. Whittle [443]. Who [91]. Wickens [605]. William [492]. Williams [418, 541, 140, 570]. Wilms [188]. Wilson [467, 491]. Windows [530]. Withdrawal [424]. Witherspoon [164, 124]. Witness [618]. Wittink [52]. Wold [203]. Women [742, 487, 145]. Woodbury [574]. Word [459]. Work [145]. Working [706, 197]. Works [720, 468]. World [585, 145, 166, 277, 351]. World-Class [277]. Wright [475]. Wynn [175, 537].

XXIII [187]. XXIX [307]. XXVI [207, 245]. XXVII [467]. XXVIII [332].

Yao [600]. Yates [426]. Year [145]. Years [214]. Young [3, 467].

Z [50, 442]. Zeger [434]. Zero [759, 729, 549]. Zero-Inflated [759, 729]. Zund [357].

References

Anonymous:1990:FMa

[1] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):??, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                             URL http://www.jstor.org/stable/
                                             2983092.

Heath:1990:CIE

[2] Anthony F. Heath and Peter Clifford. Class inequalities in education in the Twentieth Century. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):1–16, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                             www.jstor.org/stable/2983093.

Power:1990:HCS

[3] C. Power, O. Manor, A. J. Fox, and K. Fogelman. Health in childhood and social inequalities in health in young adults. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):17–28, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                             www.jstor.org/stable/2983094.

Lloyd:1990:EEA

[4] C. J. Lloyd. Estimating the effect of alcohol on the risk of a fatal road accident. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):29–52, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                             jstor.org/stable/2983095.

Smith:1990:UPI

[5] Peter Smith. The use of performance indicators in the public sector. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1): 53–72, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                             jstor.org/stable/2983096.

Kotz:1990:LSN

[6] Samuel Kotz and Eugene Seneta. Lenin as a statistician: A non-Soviet view. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153 (1):73–94, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983097.

Anonymous:1990:SSa

[7] Anonymous. Statistics in society, 1840, 1890, 1940. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):95–97, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983098.

Draper:1990:ORC

[8] Norman R. Draper. Obituary: Raj Chandra Bose. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):98–99, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983099.

Deheuvels:1990:ODD

[9] Paul Deheuvels. Obituary: Daniel Dugué. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):99–100, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983100.

David:1990:OES

[10] Florence N. David. Obituary: Elizabeth Scott, 1917–88. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):100, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983101.

Goodall:1990:BRB

[11] E. A. Goodall. Book review: Very Small Farms: An Economic Study, by D. J. Ansell; A. K. Giles; J. R. Rendell. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):101–102, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983102.

Jones:1990:BRBa

[12] P. W. Jones. Book review: Bayesian Statistics 3: Proceedings of the Third Valencia International Meeting, June 1st–5th, 1987, by J. M. Bernardo; M. H. DeGroot; D. V. Lindley; A. F. M. Smith. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):102, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983103.

Osmond:1990:BRB

[13] Clive Osmond. Book review: Statistical Methods in Cancer Research, Vol. 2, The Design and Analysis of Cohort Studies, by N. E. Breslow; N. E. Day. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153 (1):102–103, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983104.

Weatherup:1990:BRB

[14] S. T. C. Weatherup. Book review: Tables for Selecting the Best Population, by K. L. Butler; D. G. Butler. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1): 103–104, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983105.

Fielding:1990:BRB

[15] A. Fielding. Book review: Sensitivity Analysis in Linear Regression, by S. Chatterjee; A. S. Hadi. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):104, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983106.

Lovie:1990:BRBa

[16] Pat Lovie. Book review: Unified Theory and Strategies of Survey Sampling, by A. Chaudhuri; J. W. E. Vos. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1): 104–105, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983107.

Selby:1990:BRB

[17] M. J. P. Selby. Book review: The Inverse Gaussian Distribution, by R. S. Chhikara; J. L. Folks. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1): 105–106, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983108.

Phillips:1990:BRB

[18] M. J. Phillips. Book review: Accelerated Life Testing and Experts’ Opinion in Reliability, by C. A. Clarotti; D. V. Lindley. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):106–107, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983109.

Lovie:1990:BRBb

[19] Pat Lovie. Book review: Dynamic Graphics for Statistics, by W. S. Cleveland; M. E. McGill. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):107–108, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983110.

Newby:1990:BRB

[20] Martin Newby. Book review: Parameter Estimation in Reliability and Life Span Models, by A. C. Cohen; B. Jones Whitten. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):108–109, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983111.

Bailey:1990:BRBa

[21] B. J. R. Bailey. Book review: Principles of Random Variate Generation, by J. Dagpunar. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):109–110, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983112.

Metcalfe:1990:BRB

[22] A. V. Metcalfe. Book review: Quality Control, Robust Design and the Taguchi Method, by K. Dehnad. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1): 110–111, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983113.

Bailey:1990:BRBb

[23] R. A. Bailey. Book review: Comparative Experiments with Field Crops, by G. V. Dyke. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):111–112, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983114.

Milo:1990:BRB

[24] Władysław Milo. Book review: Multivariate Statistics: A Practical Approach, by B. Flury; H. Riedwyl. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1): 112–113, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983115.

Baxter:1990:BRB

[25] L. A. Baxter. Book review: A Celebration of Applied Probability, by J. Gani. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):113–114, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983116.

Stephens:1990:BRB

[26] Joan Stephens. Book review: Mastering Statistics, by T. J. Hannagan.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 153 (1):114–115, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983117.

BhrolchAin:1990:BRB

[27] Máire Ní Bhrolcháin. Book review: The Changing Population of Britain, by H. Joshi. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):115–116, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983118.

Ray:1990:BRBa

[28] W. D. Ray. Book review: Statistical Process Control in Automated Manufacturing, by J. B. Keats; N. F. Hubele. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153 (1):116–117, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983119.

Lovie:1990:BRBc

[29] Pat Lovie. Book review: Statistical Theory and Data Analysis II, by K. Matusita. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):117–118, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983120.

Stroinski:1990:BRB

[30] Krzysztof Stroiński. Book review: Taking Risks: The Science of Uncertainty, by P. Sprent. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):118, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983121.

Fayers:1990:BRB

[31] Peter Fayers. Book review: The Visual Display of Quantitative Information, by E. R. Tufte. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):118–119, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983122.

Anonymous:1990:BMa

[32] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(1):??, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983123.

Anonymous:1990:FMb

[33] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):??, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982797.

Lievesley:1990:RSS

[34] Denise Lievesley. Royal Statistical Society meeting on public confidence in the integrity and validity of official statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):121, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982798.

Hibbert:1990:PCI

[35] J. Hibbert. Public confidence in the integrity and validity of official statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):123–150, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982799.

Mead:1990:NOD

[36] R. Mead. The non-orthogonal design of experiments. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):151–201, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982800.

Rudd:1990:ECS

[37] Ernest Rudd. The early careers of social science graduates and the value of a PhD. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):203–232, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982801.

Burnhill:1990:PES

[38] Peter Burnhill, Cathy Garner, and Andrew McPherson. Parental education, social class and entry to higher education 1976–86. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):233–248, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982802.

Anonymous:1990:SSb

[39] Anonymous. Statistics in society, 1840, 1890, 1940. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):249–250, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982803.

Lindley:1990:OMD

[40] D. V. Lindley. Obituary: Morris De-Groot, 1931–89. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):251–252, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982804.

Thompson:1990:OMJ

[41] F. P. Thompson. Obituary: Michael Joseph Moroney, BSc, FSS, FIS. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2): 252–253, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982805.

Mardia:1990:OPB

[42] K. V. Mardia. Obituary: Professor B. L. Welch. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):253–254, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982806.

Chalmers:1990:BRBa

[43] Colin P. Chalmers. Book review: Causal Relationships in Medicine: A Practical System for Critical Appraisal, by J. M.

Elwood. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):255–256, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982807.

Jolliffe:1990:BRBa

[44] F. R. Jolliffe. Book review: Statistics with Confidence: Confidence Intervals and Statistical Guidelines, by M. J. Gardner; D. G. Altman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):256, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982808.

Bather:1990:BRB

[45] J. A. Bather. Book review: Multi-Armed Bandit Allocation Indices, by J. C. Gittins. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):257, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982809.

Bingham:1990:BRB

[46] N. H. Bingham. Book review: An Introduction to the Theory of Coverage Processes, by P. Hall. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2): 257–258, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982810.

Shoesmith:1990:BRB

[47] Eddie Shoesmith. Book review: Mastering the SAS System, by J. A. Jaffe.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 153 (2):258–259, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982811.

Jolliffe:1990:BRBb

[48] Ian T. Jolliffe. Book review: Simulation Methodology for Statisticians, Operations Analysts, and Engineers, Vol. 1, by P. A. W. Lewis; E. J. Orav. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153 (2):259–260, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982812.

Porter:1990:BRB

[49] M. A. Porter. Book review: What Every Engineer Should Know about Quality Control, by T. Pyzdek. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2): 260–261, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982813.

Bagley:1990:BRB

[50] J. H. Bagley. Book review: Exercise Manual in Probability Theory, by J. Stoyanov; I. Mirazchüski; Z. Ignatov; M. Tanushev. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):261, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982814.

Carroll:1990:BRB

[51] Patrick Carroll. Book review: Non-Life Insurance Mathematics, by E. Straub. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):262, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982815.

Chalmers:1990:BRBb

[52] Colin P. Chalmers. Book review: The Application of Regression Analysis, by D. R. Wittink. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):262–263, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982816.

Anonymous:1990:BMb

[53] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(2):??, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982817.

Anonymous:1990:VI

[54] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153 (3):404, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982972.

Anonymous:1990:FMc

[55] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3):??, ????

1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982973.

Moore:1990:SCN

[56] Peter G. Moore. The skills challenge of the nineties. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3):265–285, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982974.

Copas:1990:RLS

[57] J. B. Copas and F. J. Hilton. Record linkage: Statistical models for matching computer records. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3):287–320, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982975.

Bartlett:1990:CC

[58] M. S. Bartlett. Chance or chaos? Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3): 321–347, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982976.

Diggle:1990:PPM

[59] Peter J. Diggle. A point process modelling approach to raised incidence of a rare phenomenon in the vicinity of a prespecified point. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3):349–362, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982977.

Anonymous:1990:RCS

[60] Anonymous. Report of the Council for the session 1989–90. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3): 363–393, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982978.

Anonymous:1990:SSc

[61] Anonymous. Statistics in society, 1840, 1890, 1940. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3):395–396, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982979.

Nelder:1990:ODB

[62] J. A. Nelder. Obituary: Dr Brian Edward Cooper, 1935–90. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3):397, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982980.

Arber:1990:BRB

[63] Sara Arber. Book review: Population Statistics: A Review of UK Source, by B. Benjamin. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3):398–399, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982981.

McHale:1990:BRB

[64] David McHale. Book review: Pratique des Chroniques et de la Prévision à Court Terme, by A. Bensaber; B. Bleuse-Trillon. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3):399, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982982.

Ray:1990:BRBb

[65] W. D. Ray. Book review: Time Series: Theory and Methods, by P. J. Brockwell; R. A. Davis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3):400, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982983.

Jones:1990:BRBb

[66] P. W. Jones. Book review: Readings in Decision Analysis, by S. French. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3): 400–401, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982984.

Reid:1990:BRB

[67] N. G. Reid. Book review: Statistical Reasoning in Law and Public Policy, Vol. 2, Tort Law, Evidence and Health, by J. L. Gastwirth. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3): 401–402, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982985.

Farebrother:1990:BRB

[68] R. W. Farebrother. Book review: Nonlinear L_{p}-Norm Estimation, by R. Gonin; A. H. Money. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3):402, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982986.

Ripley:1990:BRB

[69] B. D. Ripley. Book review: Statistical Analysis of Random Fields, by A. V. Ivanov; N. N. Leonenko. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3): 402–403, ???? 1990. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982987.

Grimmett:1990:BRB

[70] G. R. Grimmett. Book review: General Theory of Markov Processes, by M. Sharpe. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3):403, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982988.

Anonymous:1990:BMc

[71] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153(3):??, ???? 1990. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982989.

Anonymous:1991:FMa

[72] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):??, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982686.

Gore:1991:MPC

[73] Sheila M. Gore, Tim Holt, and I. P. Fellegi. Maintaining public confidence in official statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):1–6, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982687.

Castles:1991:RUN

[74] Ian Castles. Responding to user needs. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154 (1):6–10, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982688.

Fienberg:1991:ISM

[75] Stephen E. Fienberg. Innovative statistical methodology enables ‘counting with confidence’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):10–14, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982689.

Oien:1991:OJR

[76] Arne Øien. The organization, jurisdiction and responsibilities of official statistical services. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):15–19, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982690.

Franchet:1991:ICS

[77] Yves Franchet. International comparability of statistics: Background, harmonization principles and present issues. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154 (1):19–22, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982691.

Anonymous:1991:OSC

[78] Anonymous. Official statistics: Counting with confidence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):23–44, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982692.

Aitken:1991:I

[79] C. G. G. Aitken. [introduction]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1): 45–48, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982693.

Fienberg:1991:SEU

[80] Stephen E. Fienberg and Miron L. Straf. Statistical evidence in the US courts: An appraisal. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):49–59, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982694.

Fienberg:1991:LSA

[81] Stephen E. Fienberg and D. H. Kaye. Legal and statistical aspects of some mysterious clusters. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):61–74, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982695.

Kaye:1991:CJU

[82] D. H. Kaye and Jonathan J. Koehler. Can jurors understand probabilistic evidence? Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):75–81, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982696.

Lindley:1991:SPD

[83] Dennis V. Lindley. Subjective probability, decision analysis and their legal consequences. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):83–92, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982697.

Bisgaard:1991:DSR

[84] Søren Bisgaard. Design of standards and regulations. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):93–96, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982698.

Curnow:1991:ADA

[85] R. N. Curnow and Sharon Wheeler. Assessment of deoxyribonucleic acid fingerprinting evidence in paternity and immigration cases. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):97–99, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982699.

Fatti:1991:ATP

[86] L. Paul Fatti and Michael J. Greenacre. Antitar tooth-paste: The statistical story. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):101–105, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982700.

Jackson:1991:TDT

[87] John D. Jackson. Towards a dialetic theory of proof for legal procedure. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1): 107–109, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982701.

Eisenberg:1991:RBP

[88] Theodore Eisenberg. The relationship between plaintiff success rates before trial and at trial. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):111–116, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982702.

Tarling:1991:SSC

[89] Roger Tarling. Statistics in solving crime. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):117–119, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982703.

Gastwirth:1991:PEU

[90] Joseph L. Gastwirth. The potential effect of unchecked statistical assumptions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):121–123, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982704.

Elffers:1991:URI

[91] H. Elffers, H. S. J. Robben, and D. J. Hessing. Under-reporting income: Who is the best judge–tax-payer or tax inspector? Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):125–127, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982705.

Callen:1991:CSB

[92] Craig R. Callen. Cognitive science, Bayesian norms and rules of evidence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154 (1):129–130, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982706.

Bruce:1991:ATD

[93] D. Bruce, J. R. Cuthbert, and R. E. Lygo. Assessing the teaching demands of a given school curriculum: The curriculum entry theory. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):131–141, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982707.

Besag:1991:DCR

[94] Julian Besag and James Newell. The detection of clusters in rare diseases. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154 (1):143–155, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982708.

Bhargava:1991:ESL

[95] Alok Bhargava. Estimating short and long run income elasticities of foods and nutrients for rural South India. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1): 157–174, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982709.

Anonymous:1991:SSa

[96] Anonymous. Statistics in society, 1841, 1891, 1941. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):175–177, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982710.

Kuczek:1991:OPP

[97] Thomas Kuczek. Obituary: P. S. Puri. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):178, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982711.

Parker:1991:BRB

[98] R. A. Parker. Book review: Analysis of Infectious Disease Data, by N. G.

Becker. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):179, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982712.

Jones:1991:BRB

[99] P. W. Jones. Book review: Statistical Games and Human Affairs, by R. J. Bowden. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):180, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982713.

Wickham:1991:BRB

[100] Carol Wickham. Book review: Statistics for Biologists, by R. C. Campbell. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154 (1):180–181, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982714.

Balmer:1991:BRB

[101] David Balmer. Book review: Theoretical and Computational Aspects of Simulated Annealing, by P. J. M. van Laarhoven. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154 (1):181–182, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982715.

Hasted:1991:BRB

[102] Ann Hasted. Book review: Analysis of Messy Data, Vol. 2, Non-Replicated Experiments, by G. A. Milliken; D. E.

Johnson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):182–183, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982716.

Krotki:1991:BRBa

[103] Karol J. Krotki. Book review: The Counties and Regions of the UK: Economic, Industrial and Social Trends in Local Authority Areas, by D. Mort. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154 (1):183–184, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982717.

Loukas:1991:BRB

[104] S. Loukas. Book review: Nonparametric Estimation of Probability Densities and Regression Curves, by E. A. Nadaraya. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154 (1):184–185, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982718.

Congdon:1991:BRB

[105] Peter Congdon. Book review: The Index Number Problem, by G. Stuvel. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1): 185–186, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982719.

Hitchcock:1991:BRB

[106] S. E. Hitchcock. Book review: Random Processes: A First Look, by R.

Syski. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):187, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982720.

Chatfield:1991:BRB

[107] Chris Chatfield. Book review: Empirical Model Building, by J. R. Thompson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):187–188, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982721.

Anonymous:1991:BMa

[108] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(1):??, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982722.

Anonymous:1991:FMb

[109] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):??, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983037.

Holt:1991:EC

[110] Tim Holt. Editorial: Census 1991. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2): 189–190, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983038.

Hasted:1991:SAP

[111] Anne Hasted, Roger Mead, John Sumsion, and Sharon Wheeler. Statistical analysis of public lending right loans. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154 (2):191–222, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983039.

Altman:1991:IDU

[112] Douglas G. Altman and J. Martin Bland. Improving doctors’ understanding of statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):223–267, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983040.

Ennis:1991:FFC

[113] J. R. Ennis, D. N. Wackett, and K. J. Mitchell. Fewer felons?: Crime in London in the 1990s. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):269–286, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983041.

Cole:1991:SSQ

[114] T. J. Cole, C. J. Morley, A. J. Thornton, M. A. Fowler, and P. H. Hewson. A scoring system to quantify illness in babies under 6 months of age. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154 (2):287–304, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983042.

Marsh:1991:CSA

[115] Catherine Marsh, Chris Skinner, Sara Arber, Bruce Penhale, Stan Openshaw, John Hobcraft, Denise Lievesley, and Nigel Walford. The case for samples of anonymized records from the 1991 census. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):305–340, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983043.

Newbold:1991:LEG

[116] Paul Newbold and Christos Agiakloglou.

Looking for evolving growth rates and cycles in British industrial production, 1700–1913. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):341–348, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983044.

Anonymous:1991:SSb

[117] Anonymous. Statistics in society, 1841, 1891, 1941. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):349–351, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983045.

Fellegi:1991:OMH

[118] I. P. Fellegi. Obituary: Morris H.

Hansen, 1910–90. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):352–353, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983046.

Berry:1991:OPD

[119] Geoffrey Berry. Obituary: Peter D. Oldham, 1921–89. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):353–354, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983047.

Cox:1991:OPD

[120] D. R. Cox. Obituary: Professor S. D. Silvey, 1924–90. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):354–355, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983048.

Baxter:1991:BRB

[121] L. A. Baxter. Book review: A Celebration of Statistics: The ISI Centenary Volume, by A. C. Atkinson; S. E. Fienberg. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):356–357, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983049.

Anderson:1991:BRB

[122] C. W. Anderson. Book review: Extreme Value Theory in Engineering, by E. Castillo. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):357–358, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983050.

Daly:1991:BRB

[123] Fergus Daly. Book review: Point Processes, by D. R. Cox; V. Isham. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2): 358–359, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983051.

Whitehead:1991:BRB

[124] Frank Whitehead. Book review: British Social Attitudes–Special International Report, by R. Jowell; S. Witherspoon; L. Brook. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):359–360, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983052.

Krzanowski:1991:BRB

[125] W. J. Krzanowski. Book review: Three-Way Methods for the Analysis of Qualitative and Quantitative Two-Way Data, by H. A. L. Kiers. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):360–361, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983053.

Burridge:1991:BRB

[126] Jim Burridge. Book review: Generalized Linear Models, by P. McCullagh; J. A. Nelder. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):361–364, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983054.

Krotki:1991:BRBb

[127] Karol J. Krótki. Book review: Formal Theories in International Relations, by M. Nicholson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):364–365, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983055.

Gower:1991:BRB

[128] J. C. Gower. Book review: Coordinate-Free Multivariable Statistics: An Illustrated Geometric Progression from Halmos to Gauss and Bayes, by M. Stone. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154 (2):365–366, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983056.

Anonymous:1991:BMb

[129] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(2):??, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983057.

Anonymous:1991:VI

[130] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3): 495–496, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983147.

Anonymous:1991:FMc

[131] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3):??, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983148.

Johnson:1991:EMS

[132] Anne M. Johnson. Editorial: Measuring sexual behaviour. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3):367–370, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983149.

Keiding:1991:ASI

[133] Niels Keiding. Age-specific incidence and prevalence: A statistical perspective. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3):371–412, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983150.

Moore:1991:SSD

[134] Peter Moore. Statistics and statisticians in drug regulation in the United Kingdom. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3):413–419, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983151.

Marshall:1991:RMS

[135] Roger J. Marshall. A review of methods for the statistical analysis of spatial patterns of disease. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3):421–441, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983152.

Bennett:1991:RFR

[136] J. H. Bennett. R. A. Fisher and the role of a statistical consultant. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3): 443–445, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983153.

Anonymous:1991:RCS

[137] Anonymous. Report of the Council for the session 1990–91. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3): 447–478, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983154.

Anonymous:1991:SSc

[138] Anonymous. Statistics in society, 1841, 1891, 1941. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3):479–481, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983155.

Armitage:1991:OSA

[139] Peter Armitage. Obituary: Sir Austin Bradford Hill, 1897–1991. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3): 482–484, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983156.

Cormack:1991:ODA

[140] Richard Cormack. Obituary: David Anthony Williams. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3):484–485, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983157.

Chapman:1991:BRB

[141] P. G. Chapman. Book review: The New Palgrave: Econometrics, by J. Eatwell; M. Milgate; P. Newman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3): 486–487, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983158.

Thomas:1991:BRB

[142] J. J. Thomas. Book review: The Underground Economies: Tax Evasion and Information Distortion, by E. L. Feige. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154 (3):487–490, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983159.

Davidson:1991:BRB

[143] James Davidson. Book review: Misspecification Tests in Econometrics: The Lagrange Multiplier Principle and Other Approaches, by L. G. Godfrey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3): 490–491, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983160.

Dye:1991:BRB

[144] Christopher Dye. Book review: Stage-Structured Populations: Sampling, Analysis and Simulation, by B. F. J. Manly. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3):491–492, ???? 1991. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983161.

Hinde:1991:BRB

[145] P. R. A. Hinde. Book review: Women in the World of Work: Statistical Analysis and Projections to the Year 2000, by S. Nuss; E. Denti; D. Viry. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3):492, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983162.

Metcalfe:1991:BRB

[146] Andrew V. Metcalfe. Book review: Adaptive Control and Identification: Conflict or Conflux?, by J. W. Polderman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3):493, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983163.

Darton:1991:BRB

[147] R. A. Darton. Book review: Modeling Multigroup Populations, by R. Schoen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154 (3):493–494, ???? 1991. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983164.

Anonymous:1991:BMc

[148] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 154(3):??, ???? 1991. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983165.

Anonymous:1992:FMa

[149] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1):??, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982664.

McPherson:1992:EHN

[150] Klim McPherson. Editorial: The health of the nation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1):1–3, ???? 1992. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982665.

Moore:1992:NSC

[151] Peter G. Moore. A National Statistical Commission. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1):5–28, ???? 1992. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982666.

Plackett:1992:HE

[152] Robin L. Plackett. Herschel on estimation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1):29–35, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982667.

Paterson:1992:IOA

[153] Lindsay Paterson. The influence of opportunity on aspirations among prospective university entrants from Scottish schools, 1970–88. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1):37–60, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/

2982668. See corrigendum [215].

Mills:1992:MSP

[154] Terence C. Mills and Alessandra G. Mills. Modelling the seasonal patterns in UK macroeconomic time series. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1):61–75, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982669.

Geary:1992:AMD

[155] D. N. Geary, E. Huntington, and R. J. Gilbert. Analysis of multivariate data from four dental clinical trials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1): 77–89, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982670.

Holmes:1992:SAM

[156] D. I. Holmes. A stylometric analysis of Mormon scripture and related texts.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 155 (1):91–120, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982671.

Martin:1992:COM

[157] Jean Martin and David Elliot. Creating an overall measure of severity of disability for the office of population censuses and surveys disability survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155 (1):121–140, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982672.

Sillitoe:1992:EGB

[158] K. Sillitoe and P. H. White. Ethnic group and the British Census: The search for a question. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1): 141–163, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982673.

Collins:1992:LE

[159] Martin Collins, C. Marsh, C. Skinner, S. Arber, B. Penhale, S. Openshaw, J. Hobcraft, D. Lievesley, and N. Walford. Letter to the editors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1): 165–167, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982674.

Anonymous:1992:SSa

[160] Anonymous. Statistics in society, 1842, 1892, 1942. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1):168–170, ???? 1992. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982675.

Norris:1992:BRB

[161] J. R. Norris. Book review: Disorder in Physical Systems: A Volume in Honour of John M. Hammersley, by G. R. Grimmett; D. J. A. Welsh. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1): 171–172, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982676.

Edwards:1992:BRB

[162] A. W. F. Edwards. Book review: A History of Probability and Statistics and Their Applications Before 1750, by A. Hald. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1):172–173, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982677.

Krzanowski:1992:BRB

[163] W. J. Krzanowski. Book review: A User’s Guide to Principal Components, by J. E. Jackson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1):173–174, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982678.

Siltanen:1992:BRB

[164] Janet Siltanen. Book review: British Social Attitudes–The 7th Report, by R. Jowell; S. Witherspoon; L. Brook. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1): 174–175, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982679.

Jacka:1992:BRB

[165] S. D. Jacka. Book review: Stochastic Flows and Stochastic Differential Equations, by H. Kunita. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1):175–176, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982680.

Hinde:1992:BRB

[166] P. R. A. Hinde. Book review: Expectations of Life: A Study in the Demography, Statistics, and History of World Mortality, by H. O. Lancaster. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1): 176–177, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982681.

Ford:1992:BRB

[167] Ian Ford. Book review: A Dictionary of Statistical Terms, by F. H. C. Marriott. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1):177–178, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982682.

Atkinson:1992:BRB

[168] Anthony Atkinson. Book review: Subset Selection in Regression, by A. J. Miller. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1):178–179, ???? 1992. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982683.

Grubb:1992:BRB

[169] Howard Grubb. Book review: Optimal Scaling of Time Series, by S. Van Buuren. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1):179–180, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982684.

Anonymous:1992:BMa

[170] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(1):??, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982685.

Anonymous:1992:FMb

[171] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2):??, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982952.

Smith:1992:ECS

[172] T. M. F. Smith. Editorial: The central statistical office and agency status. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155 (2):181–184, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982953.

Neumann:1992:FHD

[173] Peter M. Neumann. The future for honours degree courses in mathematics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155 (2):185–189, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982954. See dis-

cussion [175].

Wynn:1992:SDC

[174] H. P. Wynn. Statistics degree courses.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 155 (2):191–195, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982955. See dis-

cussion [175].

Anonymous:1992:DPN

[175] Anonymous. Discussion of the papers by Neumann and Wynn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2): 197–220, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982956. See [173,

174].

Bhargava:1992:MRI

[176] Alok Bhargava. Malnutrition and the role of individual variation with evidence from India and the Philippines. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155 (2):221–231, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982957.

Olaloye:1992:MCU

[177] A. O. Olaloye. Measurement of capital usage in Nigerian industries. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2): 233–239, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982958.

Cohen:1992:FPM

[178] G. Cohen and E. J. P. Fraser. Female participation in mathematical degrees at English and Scottish universities. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2):241–258, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982959.

Walker:1992:ARE

[179] Monica A. Walker. Arrest rates and ethnic minorities: A study in a provincial city. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2):259–272, ???? 1992. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982960.

Johnson:1992:THO

[180] Paul Johnson and Steven Webb. The treatment of housing in official low income statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2):273–290, ????

1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982961.

Cox:1992:CSS

[181] D. R. Cox. Causality: Some statistical aspects. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2):291–301, ???? 1992. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982962.

Anonymous:1992:SSb

[182] Anonymous. Statistics in society, 1842, 1892, 1942. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2):303–306, ???? 1992. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982963.

Prescott:1992:BRB

[183] P. Prescott. Book review: Order Statistics and Inference–Estimation Methods, by N. Balakrishnan; A. C. Cohen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2):307, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982964.

Smith:1992:BRB

[184] A. F. M. Smith. Book review: Elements of Bayesian Statistics, by J.- P. Florens; M. Mouchart; J.-M. Rolin. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2):308, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982965.

Breeze:1992:BRB

[185] Peter Breeze. Book review: Point Processes and Their Statistical Inference, by A. F. Karr. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2):308–309, ???? 1992. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982966.

Kiernan:1992:BRB

[186] Kathleen E. Kiernan. Book review: Future Demographic Trends in Europe and North America: What Can We Assume Today?, by W. Lutz. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2):309, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982967.

Hill:1992:BRB

[187] Berkeley Hill. Book review: Reviews of United Kingdom Statistical Sources: Vol. XXIII, Agriculture, by G. H. Peters. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2):310–311, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982968.

Hutchinson:1992:BRB

[188] Dougal Hutchinson. Book review: Schools, Classrooms and Pupils: International Studies of Schooling from a Multilevel Perspective, by S. W. Raudenbush; J. D. Wilms. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2):311–312, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982969.

Jackson:1992:BRB

[189] David A. Jackson. Book review: Dynamic Programming: A Practical Introduction, by D. K. Smith. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2):312, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982970.

Anonymous:1992:BMb

[190] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(2):??, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982971.

Anonymous:1992:VI

[191] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3): 479–480, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982885.

Anonymous:1992:FMc

[192] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):??, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982886.

Goldstein:1992:ESI

[193] Harvey Goldstein. Editorial: Statistical information and the measurement of educational outcomes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3): 313–315, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982887.

Thunhurst:1992:MHU

[194] Colin Thunhurst and Alison MacFarlane. Monitoring the health of urban populations: What statistics do we need? Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):317–352, ???? 1992. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982888.

Cox:1992:QLA

[195] D. R. Cox, R. Fitzpatrick, A. E. Fletcher, S. M. Gore, D. J. Spiegelhalter, and D. R. Jones. Quality-of-life assessment: Can we keep it simple? Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):353–393, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982889.

Upton:1992:FET

[196] Graham J. G. Upton. Fisher’s exact test. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):395–402, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982890.

Sheldon:1992:RAR

[197] Trevor A. Sheldon and Roy Carr-Hill. Resource allocation by regression in the National Health Service: A critique of the Resource Allocation Working Party’s review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):403–420, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982891.

Anonymous:1992:RCS

[198] Anonymous. Report of the Council for the session 1991–92. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3): 421–455, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982892.

Mills:1992:LE

[199] T. C. Mills, P. Newbold, and C. Agiakloglou. Letter to the editors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3): 457–459, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982893.

Anonymous:1992:SSc

[200] Anonymous. Statistics in society, 1842, 1892, 1942. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):460–463, ???? 1992. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982894.

Freedman:1992:ODP

[201] Laurence S. Freedman. Obituary:

David Peery Byar (1938–91). Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):

464–465, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982895.

Coutts:1992:OSB

[202] Stewart Coutts. Obituary: Sidney Benjamin. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):465–466, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982896.

Whittle:1992:OPH

[203] P. Whittle. Obituary: Professor Herman Wold. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):466–469, ???? 1992. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982897.

Weale:1992:OSR

[204] M. R. Weale. Obituary: Sir Richard Stone. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):469–470, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982898.

Ray:1992:BRBa

[205] W. D. Ray. Book review: Statistical Analysis of Reliability and Life Testing Models, Theory and Methods, by L. J. Bain; M. Engelhardt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):471, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982899.

Skinner:1992:BRBa

[206] C. J. Skinner. Book review: Sample Survey Principles and Methods, by V. Barnett. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):472, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982900.

Clegg:1992:BRB

[207] Jeremy Clegg. Book review: Reviews of United Kingdom Statistical Sources: Vol. XXVI, International Aspects of UK Economic Activities, by P. J. Buckley; R. D. Pearce. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):472–473, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/

2982901. See corrigendum [245].

Bowman:1992:BRB

[208] Adrian Bowman. Book review: Aspects of Nonparametric Density Estimation, by A. J. van Es. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3): 473–474, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982902.

Biggins:1992:BRB

[209] J. D. Biggins. Book review: Statistical Inference for Branching Processes, by P. Guttorp. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):474, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982903.

Ray:1992:BRBb

[210] W. D. Ray. Book review: Applied Multivariate Data Analysis: Vol. 1, Regression and Experimental Design, by J. D. Jobson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):474–475, ???? 1992. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982904.

Skinner:1992:BRBb

[211] C. J. Skinner. Book review: Drawing General Population Samples in Great Britain, by P. Lynn; D. Lievesley. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):475, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982905.

Mills:1992:BRB

[212] Terence C. Mills. Book review: The Art of Statistical Science: A Tribute to G. S. Watson, by K. V. Mardia. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155 (3):475–476, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982906.

Newton:1992:BRB

[213] J. Newton. Book review: Stochastic Models in Queueing Theory, by J. Medhi. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):476–477, ???? 1992. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982907.

McClean:1992:BRB

[214] Sally McClean. Book review: Health and Class: The Early Years, by C. Power; O. Manor; J. Fox. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3): 477–478, ???? 1992. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982908.

Paterson:1992:CIO

[215] Lindsay Paterson. Corrigendum: The influence of opportunity on aspirations among prospective university entrants from Scottish schools, 1970–88. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):481, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/

2982909. See [153].

Anonymous:1992:BMc

[216] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 155(3):??, ???? 1992. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982910.

Anonymous:1993:FMa

[217] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):??, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982855.

Maclean:1993:ECC

[218] Ian Maclean. Editorial: Changes in the collection of trade statistics following the introduction of the single market in January 1993. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):1–4, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982856.

McLennan:1993:ECS

[219] W. McLennan. Editorial: Clarifying the scope of government economic statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):4–7, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982857.

Draper:1993:EDA

[220] David Draper, James S. Hodges, Colin L. Mallows, and Daryl Pregibon. Exchangeability and data analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):9–37, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982858.

Barnett:1993:EAP

[221] V. Barnett and S. Hilditch. The effect of an artificial pitch surface on home team performance in football (soccer). Journal of the Royal Statistical Society. Series A (Statistics in Society), 156 (1):39–50, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982859.

Clifford:1993:PCS

[222] P. Clifford and A. F. Heath. The political consequences of social mobility. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156 (1):51–61, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982860.

Dunn:1993:SND

[223] G. Dunn, P. C. Sham, and D. J. Hand.

Statistics and the nature of depression. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156 (1):63–87, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982861.

Solomou:1993:BEN

[224] Solomos Solomou and Martin Weale.

Balanced estimates of national accounts when measurement errors are autocorrelated: The UK, 1920–38. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1): 89–105, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982862.

Copas:1993:FAR

[225] J. B. Copas. A formula for the allocation of resources based on uncertain predictions of need. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):107–113, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982863.

Schaalje:1993:LE

[226] G. Bruce Schaalje and D. I. Holmes. Letter to the editors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):115–117, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982864.

Anonymous:1993:SSa

[227] Anonymous. Statistics in society, 1843, 1893, 1943. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):118–120, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982865.

Redfern:1993:OGT

[228] Philip Redfern. Obituary: George (‘Toby’) Paine CB, DFC, 1918–92. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1): 121–122, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982866.

Freeman:1993:BRBa

[229] Jim Freeman. Book review: Statistics Tutor: Tutorial and Computational Software for the Behavioral Sciences, by J. D. Allen; D. J. Pittenger. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1): 123–124, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982867.

Green:1993:BRBa

[230] M. W. Green. Book review: Handbook of the Logistic Distribution, by N. Balakrishnan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):124, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982868.

Ross:1993:BRB

[231] Gavin J. S. Ross. Book review: Correspondence Analysis Handbook, by J.-P. Benzecri. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):124–125, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982869.

Brigden:1993:BRB

[232] M. E. Brigden. Book review: The Linear Complementarity Problem, by R. W. Cottle; J.-S. Pang; R. E. Stone. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1): 125–126, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982870.

Rees:1993:BRB

[233] Mike Rees. Book review: Statistics for Spatial Data, by N. Cressie. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1): 126–127, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982871.

Humphries:1993:BRB

[234] Jane Humphries. Book review: Lone Parenthood: An Economic Analysis, by J. F. Ermisch. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):127–128, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982872.

Barry:1993:BRB

[235] Daniel Barry. Book review: Applied Nonparametric Regression, by W. Härdle. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):128–129, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982873.

Shoesmith:1993:BRB

[236] Eddie Shoesmith. Book review: Design and Inference in Finite Population Sampling, by A. S. Hedayat; B. K. Sinha. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156 (1):129–130, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982874.

Goldstein:1993:BRB

[237] Michael Goldstein. Book review: Scientific Reasoning: The Bayesian Approach, by C. Howson; P. Urbach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1): 130–131, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982875.

Jolliffe:1993:BRB

[238] Ian T. Jolliffe. Book review: Exploratory and Multivariate Data Analysis, by M. Jambu. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):131–132, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982876.

Davison:1993:BRB

[239] A. C. Davison. Book review: Exploring the Limits of Bootstrap, by R. LePage; L. Billard. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):133, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982877.

Fielding:1993:BRB

[240] A. Fielding. Book review: Sampling of Populations: Methods and Applications, by P. S. Levy; S. Lemeshow. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):134, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982878.

OBrien:1993:BRB

[241] Carl M. O’Brien. Book review:

Computer-Interactive Data Analysis, by A. D. Lunn; D. R. McNeil. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1): 134–135, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982879.

Freeman:1993:BRBb

[242] G. H. Freeman. Book review: Sample Size Choice, by R. E. Odeh; M. Fox. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):135–136, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982880.

Jennings:1993:BRB

[243] Simon Jennings. Book review: Biopharmaceutical Statistics for Drug Development, by K. E. Peace. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1): 136–137, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982881.

Lee:1993:BRB

[244] P. M. Lee. Book review: L’Analyse Statistique Bayesienne, by C. Robert. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156 (1):137–138, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982882.

Buckley:1993:CRU

[245] P. J. Buckley and R. D. Pearce. Corrigendum: Reviews of United Kingdom statistical sources: Vol. XXVI, international aspects of UK economic activities. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):138, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982883. See [207].

Anonymous:1993:BMa

[246] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(1):??, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982884.

Anonymous:1993:FMb

[247] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):??, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982723.

Anonymous:1993:ESQ

[248] Anonymous. Editorial: Statistics and quality. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):139–141, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982724.

Stone:1993:EKC

[249] Mervyn Stone. Editorial: Kipling on 19th-Century quality control in India. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156 (2):141–143, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982725.

Smith:1993:PSL

[250] T. M. F. Smith. Populations and selection: Limitations of statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):

144–166, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982726.

Ehrenberg:1993:PP

[251] A. S. C. Ehrenberg and J. A. Bound. Predictability and prediction. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2): 167–206, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982727.

Curnow:1993:PID

[252] R. N. Curnow and Sharon Wheeler. Probabilities of incorrect decisions in paternity cases using multilocus deoxyribonucleic acid probes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2): 207–223, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982728. See corri-

gendum [312].

Duffy:1993:ACC

[253] John C. Duffy. Alcohol consumption and control policy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):225–230, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982729.

Lemmens:1993:RDA

[254] Paul H. H. M. Lemmens. Regularity in the distribution of alcohol consumption. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):231–235, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982730.

Congdon:1993:SGL

[255] Peter Congdon. Statistical graduation in local demographic analysis and projection. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):237–270, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982731.

Johnes:1993:RPU

[256] Jill Johnes, Jim Taylor, and Brian Francis. The research performance of UK universities: A statistical analysis of the results of the 1989 research selectivity exercise. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):271–286, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982732.

Greenfield:1993:CS

[257] Tony Greenfield. Communicating statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):287–297, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982733.

Dewey:1993:LE

[258] Michael Dewey. Letter to the editors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):299, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982734.

Anonymous:1993:CRS

[259] Anonymous. Conference report: Statistics and the law. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):301–309, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982735.

Anonymous:1993:SSb

[260] Anonymous. Statistics in society, 1843, 1893, 1943. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):311–314, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982736.

Kruskal:1993:KB

[261] William Kruskal. K. A. Brownlee, 1918– 90. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):315, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982737.

Grebenik:1993:JBP

[262] E. Grebenik. Jean Bourgeois-Pichat, 1912–90. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):316, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982738.

Whitehead:1993:AMA

[263] Frank Whitehead. Abraham Manie Adelstein, 1916–92. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):316–318, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982739.

Smith:1993:BRB

[264] J. Q. Smith. Book review: Bayesian Statistics 4, by J. M. Bernardo; J. O. Berger; A. P. Dawid; A. F. M. Smith. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156 (2):319–320, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982740.

Grieve:1993:BRB

[265] A. P. Grieve. Book review: Confidence Intervals on Variance Components, by R. K. Burdick; F. A. Graybill. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2): 320–321, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982741.

Schimek:1993:BRB

[266] M. G. Schimek. Book review: Truncated and Censored Samples: Theory and Applications, by A. C. Cohen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):321, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982742.

Collett:1993:BRB

[267] Dave Collett. Book review: Alcohol and Illness: The Epidemiological Viewpoint, by J. C. Duffy. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2): 321–322, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982743.

Oliver:1993:BRB

[268] F. R. Oliver. Book review: Essentials of Econometrics, by D. Gujarati. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156 (2):322–323, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982744.

Iles:1993:BRB

[269] Terence Iles. Book review: Fundamentals of Explaratory Analysis of Variance, by D. C. Hoaglin; F. Mosteller; J. W. Tukey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):323–324, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2982745.

ORegan:1993:BRB

[270] Myra O’Regan. Book review: Panel Surveys, by D. Kasprzyk; G. Duncan; G. Kalton; M. P. Singh. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):324, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982746.

Hirst:1993:BRB

[271] David Hirst. Book review: Fundamentals of Biostatistical Inference, by C. T.

Le. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156 (2):324–325, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982747.

Mills:1993:BRB

[272] Terence C. Mills. Book review: Introduction to Multiple Time Series Analysis, by H. Lutkepohl. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):325–326, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982748.

Harding:1993:BRB

[273] E. F. Harding. Book review: Quadratic Forms in Random Variables: Theory and Applications, by A. M. Matthai; S. B. Provost. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):326–327, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2982749.

Ford:1993:BRB

[274] Ian Ford. Book review: Discriminant Analysis and Statistical Pattern Recognition, by G. J. McLachlan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2): 327–328, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2982750.

Thompson:1993:BRB

[275] J. W. Thompson. Book review: Statistical Methods for the Process Industries, by W. H. McNeese; R. A. Klein.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):328, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982751.

Whitehead:1993:BRB

[276] John Whitehead. Book review: Biopharmaceutical Sequential Statistical Applications, by K. E. Peace. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):329, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2982752.

Newton:1993:BRB

[277] David Newton. Book review: Quality Engineering Handbook, by T. Pyzdek; R. W. Berger; Managing for World-Class Quality — A Primer for Executives and Managers by E. S. Shecter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156 (2):329–331, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982753.

Armitage:1993:BRB

[278] Peter Armitage. Book review: The Design and Analysis of Sequential Clinical Trials, 2nd edn, by J. Whitehead. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156 (2):331–332, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2982754.

Anonymous:1993:BMb

[279] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(2):??, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2982755.

Anonymous:1993:VI

[280] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3): 515–518, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983058.

Anonymous:1993:FMc

[281] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):??, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983059.

Derbyshire:1993:ECL

[282] Mike Derbyshire. Editorial: Changes in local government finance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3): 333–335, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983060.

Walker:1993:EGS

[283] John Walker. Editorial: Government statistical service views on the European Community Draft Data Protection Directive. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):335–338, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983061.

Bailey:1993:LTD

[284] R. E. Bailey and M. J. Chambers. Longterm demographic interactions in precensus England. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):339–362, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983062.

Lawson:1993:AME

[285] Andrew B. Lawson. On the analysis of mortality events associated with a prespecified fixed point. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3): 363–377, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983063.

Kronmal:1993:SCF

[286] Richard A. Kronmal. Spurious correlation and the fallacy of the ratio standard revisited. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):379–392, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983064.

Raboud:1993:ERP

[287] J. Raboud, N. Reid, R. A. Coates, and V. T. Farewell. Estimating risks of progressing to aids when covariates are measured with error. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):

393–406, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983065.

Wadsworth:1993:MNS

[288] J. Wadsworth, J. Field, A. M. Johnson, S. Bradshaw, and K. Wellings. Methodology of the national survey of sexual attitudes and lifestyles. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3): 407–421, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983066.

Robertson:1993:PPC

[289] Chris Robertson. Participation in postcompulsory education in Scotland. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3): 423–442, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983067.

Kimber:1993:SAB

[290] Alan C. Kimber and Alan R. Hansford. A statistical analysis of batting in cricket. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):443–455, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983068.

Anonymous:1993:RCS

[291] Anonymous. Report of the Council for the session 1992–93. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):

457–493, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983069.

Anonymous:1993:SSc

[292] Anonymous. Statistics in society, 1843, 1893, 1943. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):495–497, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983070.

Osmond:1993:PMG

[293] Clive Osmond. Professor Martin Gardner (1940–93). Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):498–499, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983071.

Goldfield:1993:RE

[294] Edwin D. Goldfield and Marie D. Wann. A. Ross Eckler. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):499–500, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983072.

Prescott:1993:BRB

[295] P. Prescott. Book review: A First Course in Order Statistics, by B. C. Arnold; N. Balakrishnan; H. N. Nagaraja. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):501–502, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983073.

Osmond:1993:BRB

[296] Clive Osmond. Book review: Applied Multivariate Data Analysis, by B. S. Everitt; G. Dunn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):502, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983074.

Jones:1993:BRB

[297] M. C. Jones. Book review: Nonparametric Statistical Inference, 3rd edn, by J. D. Gibbons; S. Chakraborti. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):503, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983075.

Doney:1993:BRB

[298] R. A. Doney. Book review: Probability and Random Processes, 2nd edn, by G. R. Grimmett; D. R. Stirzaker. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156 (3):503–504, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983076.

Ray:1993:BRBa

[299] W. D. Ray. Book review: Probability and Random Processes: Problems and Solutions, by G. R. Grimmett; D. R. Stirzaker. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):504, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983077.

Young:1993:BRB

[300] G. Alastair Young. Book review: The Bootstrap and Edgeworth Expansion, by P. Hall. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):504–505, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983078.

Omar:1993:BRB

[301] Rumana Omar. Book review: Multiple Comparisons, Selection Procedures and Applications in Biometry, by F. M. Hoppe. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):505–506, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983079.

Ray:1993:BRBb

[302] W. D. Ray. Book review: Applied Multivariate Darta Analysis: Vol. 2, Categorical and Multivariate Methods, by J. D. Jobson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):506, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983080.

Crowder:1993:BRB

[303] Martin Crowder. Book review: Point Processes and Their Statistical Inference, 2nd edn, by A. F. Karr. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):507, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983081.

Alexander:1993:BRB

[304] C. O. Alexander. Book review: Leading Economic Indicators–New Approaches and Forecasting Records, by K. Lahiri; G. H. Moore. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):507–508, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983082.

Wolff:1993:BRB

[305] Rodney C. L. Wolff. Book review: Fractal Attraction: A Fractal Design System for the Macintosh, by K. D. Lee; Y. Cohen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):508–509, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983083.

Kelly:1993:BRB

[306] Peter Kelly. Book review: Algorithms, Routines, and S Functions for Robust Statistics, by A. Marazzi. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3): 509–510, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983084.

Davies:1993:BRB

[307] Gary Davies. Book review: Reviews of United Kingdom Statistical Sources: Vol. XXIX, Distribution, by C. Moir; J. A. Dawson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):510–511, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983085.

Coad:1993:BRB

[308] D. S. Coad. Book review: Optimal Sequential Planned Decision Procedures, by N. Schmitz. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):511, ???? 1993. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983086.

Nason:1993:BRB

[309] Guy Nason. Book review: Multivariate Density Estimation: Theory, Practice, and Visualization, by D. W. Scott. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156 (3):511–512, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983087.

Green:1993:BRBb

[310] J. R. Green. Book review: Cross-Over Trials in Clinical Research, by S. Senn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156 (3):512–513, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983088.

Skinner:1993:BRB

[311] C. J. Skinner. Book review: Sampling, by S. K. Thompson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3): 513–514, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983089.

Curnow:1993:CPI

[312] R. N. Curnow and Sharon Wheeler.

Corrigendum: Probabilities of incorrect decisions in paternity cases using multilocus deoxyribonucleic acid probes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):518, ???? 1993. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983090. See [252].

Anonymous:1993:BMc

[313] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 156(3):??, ???? 1993. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983091.

Anonymous:1994:FMa

[314] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):??, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983499.

Thompson:1994:EN

[315] Simon Thompson and Tim Holt. Editorial note. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):1, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983500.

Aalen:1994:UHI

[316] O. O. Aalen, V. T. Farewell, D. De Angelis, and N. E. Day. The use of human immunodeficiency virus diagnosis information in monitoring the acquired immune deficiency syndrome epidemic. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):3–16, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983501.

Raab:1994:BFH

[317] Gillian M. Raab, Sheila M. Gore, David J. Goldberg, and Christl A. Donnelly. Bayesian forecasting of the human immunodeficiency virus epidemic in Scotland. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):17–30, ???? 1994. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983502.

DeAngelis:1994:EAI

[318] D. De Angelis and W. R. Gilks. Estimating acquired immune deficiency syndrome incidence accounting for reporting delay. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):31–40, ???? 1994. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983503.

Sabin:1994:UBE

[319] Caroline A. Sabin, Christine A. Lee, and Andrew N. Phillips. The use of backcalculation to estimate the prevalence of severe immunodeficiency induced by human immunodeficiency virus in England and Wales. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):41–56, ???? 1994. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983504.

Gore:1994:NRA

[320] Sheila M. Gore, Barry G. Evans, Daniela de Angelis, and Roland Salmon. A note on regional acquired immune deficiency syndrome projections. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1): 57–68, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983505.

Williams:1994:MMT

[321] John R. Williams and Roy M. Anderson. Mathematical models of the transmission dynamics of human immunodeficiency virus in England and Wales: Mixing between different risk groups. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157 (1):69–87, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983506.

Giesecke:1994:EPH

[322] Johan Giesecke, Anne Johnson, Anne Hawkins, Ahilya Noone, and Angus Nicoll. An estimate of the prevalence of human immunodeficiency virus infection in England and Wales by using a direct method. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):89–103, ???? 1994. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983507.

Evans:1994:CRC

[323] B. G. Evans and A. McCormick. Completeness of reporting cases of acquired immune deficiency syndrome by clinicians. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):105–114, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983508.

Mollison:1994:EMD

[324] Denis Mollison, Valerie Isham, and Bryan Grenfell. Epidemics: Models and data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):115–149, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983509.

Haigh:1994:PH

[325] John Haigh. Philip Holgate 1934–93. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157 (1):151–153, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983510.

Sprent:1994:JGP

[326] P. Sprent. E. J. G. Pitman, 1897–1993. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157 (1):153–154, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983511.

Hardy:1994:BRB

[327] Mary Hardy. Book review: Mortality on the Move, by B. Benjamin; A. S. Soliman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):155–156, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983512.

Bhrolchain:1994:BRB

[328] Maire Ni Bhrolchain. Book review: The British Population: Patterns, Trends, and Processes, by D. Coleman; J. Salt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157 (1):156–157, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983513.

Aitken:1994:BRB

[329] C. G. G. Aitken. Book review: Likelihood, by A. W. F. Edwards. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1): 157–158, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983514.

Ray:1994:BRBa

[330] W. D. Ray. Book review: Longitudinal Data with Serial Correlation: A State Space Approach, by R. H. Jones. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157 (1):158–159, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983515.

Isham:1994:BRB

[331] Valerie Isham. Book review: Sexual Behaviour and AIDS in Britain, by E. G. Knox; C. MacArthur; K. J. Simons.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 157 (1):159–160, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983516.

Lund:1994:BRB

[332] P. J. Lund. Book review: Reviews of United Kingdom Statistical Sources: Vol XXVIII, The Food Industries, by J. Mark; R. Strange. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):160–162, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983517.

Moore:1994:BRB

[333] B. A. Moore. Book review: Discriminant Analysis and Statistical Pattern Recognition, by G. J. McLachlan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):162, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983518.

Burrell:1994:BRB

[334] Quentin L. Burrell. Book review: Identifiability in Stochastic Models: Characterization of Probability Distributions, by B. L. S. Prakasa Rao. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1): 163–164, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983519.

Smith:1994:BRB

[335] J. Q. Smith. Book review: Markov Decision Processes, by D. J. White. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):164, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983520.

Elliott:1994:SIO

[336] Caroline Elliott and Richard William Farebrother. Subject index of obituary notices 1839–1993. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):165–171, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983521.

Anonymous:1994:BMa

[337] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(1):??, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983522.

Anonymous:1994:FMb

[338] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2):??, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983355.

Tarling:1994:EEI

[339] Roger Tarling. Editorial: The effect of imprisonment on crime. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2): 173–176, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983356.

Cuthbert:1994:EIE

[340] J. R. Cuthbert. An extension of the induced exposure method of estimating driver risk. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2):177–190, ???? 1994. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983357.

Campbell:1994:TSR

[341] M. J. Campbell. Time series regression for counts: An investigation into the relationship between sudden infant death syndrome and environmental temperature. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2):191–208, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983358.

Box:1994:SQI

[342] George Box. Statistics and quality improvement. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2):209–229, ???? 1994. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983359.

Upton:1994:PBG

[343] Graham J. G. Upton. Picturing the 1992 British General Election. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2):

231–252, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983360.

Gregg:1994:CSS

[344] Paul Gregg. Out for the count: A social scientist’s analysis of unemployment statistics in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2): 253–270, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983361.

Raab:1994:SCC

[345] Gillian M. Raab. Selecting confounders from covariates. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2):271–283, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983362.

Lawson:1994:ACS

[346] Andrew B. Lawson and Fiona L. R.

Williams. Armadale: A case-study in environmental epidemiology. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2): 285–298, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983363.

Barton:1994:OFN

[347] David Barton. Obituary: Florence Nightingale David. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2):299–301, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983364.

Dale:1994:OPC

[348] Angela Dale. Obituary: Professor Catherine Marsh, 1952–93. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2): 301–302, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983365.

Cuthbert:1994:ODS

[349] J. R. Cuthbert. Obituary: David Salmond. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2):302, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983366.

Brown:1994:OMJ

[350] Audrey Brown. Obituary: Miss Jean Helen Thompson, CBE. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2): 302–303, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983367.

Fielding:1994:BRB

[351] A. Fielding. Book review: Social Research on Developing Countries: Surveys and Censuses in the Third World, by M. Bulmer; D. P. Warwick. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2): 305–306, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983368.

Rudolfer:1994:BRB

[352] Stephan M. Rudolfer. Book review: Biostatistics: A Methodology for the Health Sciences, by L. D. Fisher; G. van Belle. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157 (2):306–307, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983369.

Crowder:1994:BRB

[353] Martin Crowder. Book review: Counting Processes and Survival Analysis, by T. R. Fleming; D. P. Harrington. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2): 307–308, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983370.

Ray:1994:BRBb

[354] W. D. Ray. Book review: Time Series, by Maurice Kendall; J. K. Ord. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157 (2):308–309, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983371.

Kendall:1994:BRB

[355] David Kendall. Book review: Poisson Processes, by J. F. C. Kingman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2):309, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983372.

Balmer:1994:BRB

[356] David Balmer. Book review: Simulation — A Statistical Perspective, by J. P. C. Kleijnen; W. van Groenendaal. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157 (2):309–310, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983373.

Grieve:1994:BRB

[357] A. P. Grieve. Book review: Statistical Methods in Analytical Chemistry, by P. C. Meier; R. E. Zund. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2):311, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983374.

Anonymous:1994:BMb

[358] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(2):??, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983375.

Anonymous:1994:VI

[359] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3): 512–514, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983523.

Anonymous:1994:FMc

[360] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):??, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983524.

Simpson:1994:ECG

[361] Stephen N. Simpson. Editorial: Coverage of the Great Britain census of population and housing. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):313–316, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983525.

Hand:1994:DSQ

[362] David J. Hand. Deconstructing statistical questions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):317–356, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983526.

Spiegelhalter:1994:BAR

[363] David J. Spiegelhalter, Laurence S.

Freedman, and Mahesh K. B. Parmar. Bayesian approaches to randomized trials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157 (3):357–416, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983527.

Bhargava:1994:MHF

[364] Alok Bhargava. Modelling the health of Filipino children. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):417–432, ????

1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983528.

Diggle:1994:CAP

[365] Peter J. Diggle and Barry S. Rowlingson. A conditional approach to point process modelling of elevated risk. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3): 433–440, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983529.

Islam:1994:MSM

[366] M. Ataharul Islam. Multistate survival models for transitions and reverse transitions: An application to contraceptive use data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):441–455, ???? 1994. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983530.

Anonymous:1994:RCS

[367] Anonymous. Report of the Council for the session 1993–94. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3): 457–496, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983531.

Hunter:1994:OWE

[368] J. Stuart Hunter. Obituary: W. Edwards Deming (1900–93). Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3): 497–499, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983532.

Gomes:1994:OJT

[369] M. Ivette Gomes. Obituary: J. Tiago de Oliveira. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):499–500, ???? 1994. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983533.

Keen:1994:OHC

[370] Joan Keen. Obituary: Hugo Christian Hamaker (1905–93). Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3): 500–501, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983534.

Bartholomew:1994:ORB

[371] David Bartholomew. Obituary: Robert E. Bohrer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):501–502, ???? 1994. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983535.

Jones:1994:BRB

[372] David R. Jones. Book review: Drug Safety Assessment in Clinical Trials, by G. S. Gilbert. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):503, ???? 1994. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983536.

Garthwaite:1994:BRB

[373] Paul Garthwaite. Book review: Introduction to Probability and Statistics, by N. C. Giri. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):504, ???? 1994. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983537.

Worton:1994:BRB

[374] B. J. Worton. Book review: Computer Intensive Statistical Methods: Validation Model Selection and Bootstrap, by J. S. U. Hjorth. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):504–505, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983538.

Fearn:1994:BRB

[375] Tom Fearn. Book review: Quality Control in Analytical Chemistry, by G. Kateman; L. Buydens. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):505–506, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983539.

Kimber:1994:BRB

[376] Alan Kimber. Book review: Advanced Calculus with Applications in Statistics, by A. I. Khuri. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):506, ???? 1994. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983540.

Nelder:1994:BRB

[377] J. A. Nelder. Book review: Random Coefficient Models, by N. T. Longford. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):507, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983541.

Walden:1994:BRB

[378] Andrew Walden. Book review: The Econometric Modelling of Financial Time Series, by T. C. Mills. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):508, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983542.

Roger:1994:BRB

[379] James Roger. Book review: Genstat 5 Release 3 Reference Manual, by R. W. Payne. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):508–509, ???? 1994. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983543.

Hand:1994:BRB

[380] David J. Hand. Book review: Applied Factor Analysis in the Natural Sciences, by R. Reyment; K. G. Joreskog. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):509, ???? 1994. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983544.

Mills:1994:BRB

[381] Terence C. Mills. Book review: Bounded Rationality in Macroeconomics, by T. J. Sargent. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):510, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983545.

Coleman:1994:BRB

[382] Shirley Y. Coleman. Book review: Quality: Measuring and Monitoring: (a) Lampeter and Son; (b) Studio Discussion and Case Studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):511, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983546.

Anonymous:1994:BMc

[383] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 157(3):??, ???? 1994. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983547.

Anonymous:1995:FMa

[384] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):??, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983400.

Bartholomew:1995:WS

[385] David J. Bartholomew. What is statistics? Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):1–20, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983401.

Balding:1995:IFI

[386] David J. Balding and Peter Donnelly. Inference in forensic identification. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1): 21–53, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983402.

Cleave:1995:EME

[387] N. Cleave, P. J. Brown, and C. D. Payne.

Evaluation of methods for ecological inference. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):55–72, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983403.

Rodriguez:1995:AEP

[388] German Rodriguez and Noreen Goldman. An assessment of estimation procedures for multilevel models with binary responses. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):73–89, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983404.

Harrop:1995:TDF

[389] Anne Harrop and Ian Plewis. Two decades of family change: Secondary analysis of continuous government surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):91–106, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983405.

Goetghebeur:1995:DEJ

[390] Els J. T. Goetghebeur and Stuart J. Pocock. Detection and estimation of Jshaped risk-response relationship. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1): 107–121, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983406.

Koop:1995:EIW

[391] Gary Koop and Dale J. Poirier. An empirical investigation of Wagner’s hypothesis by using a model occurrence framework. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):123–141, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983407.

Armitage:1995:BAB

[392] Peter Armitage. Before and after Bradford Hill: Some trends in medical statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):143–153, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983408.

Doll:1995:SAB

[393] Richard Doll. Sir Austin Bradford Hill: A personal view of his contribution to epidemiology. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):155–163, ????

1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983409.

Davidson:1995:OLS

[394] Roger Davidson. Official labour statistics: A historical perspective. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1): 165–173, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983410.

Goldstein:1995:GPC

[395] Harvey Goldstein and Michael J. R.

Healy. The graphical presentation of a collection of means. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):175–177, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983411.

Paddle:1995:OAB

[396] G. M. Paddle. Obituary: Albert Baines, 1927–94. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):179–180, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983412.

Boreham:1995:OGP

[397] John Boreham. Obituary: Geoffrey Penrice, 1923–94. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):180–181, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983413.

Robinson:1995:OJH

[398] P. M. Robinson. Obituary: E. J. Hannan, 1921–94. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):181–182, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983414.

Erritt:1995:OSJ

[399] John Erritt. Obituary: Sir John Boreham, KCB, 1925–94. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1): 182–184, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983415.

Peach:1995:BRB

[400] Nigel Peach. Book review: Detection of Abrupt Changes: Theory and Applications, by M. Basseville; I. V. Nikiforov. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):185–186, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983416.

Newton:1995:BRB

[401] J. Newton. Book review: Sojourn Times in Feedback and Processor Sharing Queues, by J. L. van den Berg. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158 (1):186–187, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983417.

Atkinson:1995:BRBa

[402] Anthony Atkinson. Book review: Measurement, Regression, and Calibration, by P. J. Brown. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):187, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983418.

Ray:1995:BRB

[403] W. D. Ray. Book review: Modelling Survival Data in Medical Research, by D. Collett. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):188, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983419.

Dorling:1995:BRB

[404] Daniel Dorling. Book review: The 1991 Census User’s Guide, by A. Dale; C. Marsh. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):188–189, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983420.

Ripley:1995:BRBa

[405] B. D. Ripley. Book review: Number-Theoretic Methods in Statistics, by K.- T. Fang; Y. Wang. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):189–190, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983421.

Prentice:1995:BRB

[406] M. J. Prentice. Book review: Statistical Analysis of Circular Data, by N. I. Fisher. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):190, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983422.

Dunsmore:1995:BRB

[407] Ian R. Dunsmore. Book review: Predictive Inference: An Introduction, by S. Geisser. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):191, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983423.

Davies:1995:BRBa

[408] Neville Davies. Book review: Time Series Models, 2nd edn, by A. C. Harvey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):191–193, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983424.

Davies:1995:BRBb

[409] Neville Davies. Book review: Time Series–Forecasting, Simulation, Applications, by G. Janacek; L. Swift. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1): 193–195, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983425.

Snell:1995:BRBa

[410] Joyce Snell. Book review: The Theory of Linear Models, by B. Jorgensen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):195, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983426.

Conniffe:1995:BRB

[411] D. Conniffe. Book review: Design of Experiments: A No-Name Approach, by T. J. Lorenzen; V. L. Anderson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1): 195–197, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983427.

Prescott:1995:BRB

[412] P. Prescott. Book review: Algebraic Models for Social Networks, by P. Pattison. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):197–198, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983428.

Swift:1995:BRB

[413] A. Louise Swift. Book review: Spectral Analysis for Physical Applications: Multipaper and Conventional Techniques, by D. A. Percival; A. T. Walden. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1): 198–199, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983429.

Reese:1995:BRB

[414] R. Allan Reese. Book review: Statistical Analysis of Software Reliability Models, by M. C. J. van Pul. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):199, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983430.

Walker:1995:BRBa

[415] Stephen Walker. Book review: Cross-Over Experiments: Design, Analysis, and Application, by D. A. Ratkowsky; M. A. Evans; J. R. Alldredge. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):200, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983431.

Fearn:1995:BRB

[416] Tom Fearn. Book review: Statistical Models for Causal Analysis, by R. D. Retherford; M. K. Choe. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1): 200–201, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983432.

Jolliffe:1995:BRB

[417] Flavia Jolliffe. Book review: Introduction to Probability Models, by S. M. Ross. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):201–202, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983433.

Appa:1995:BRB

[418] Gautam Appa. Book review: Model Solving in Mathematical Programming, by H. P. Williams. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):202, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983434.

Anonymous:1995:BMa

[419] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(1):??, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983435.

Anonymous:1995:FMb

[420] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):??, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983289.

Thompson:1995:CPE

[421] Eric J. Thompson. The 1991 census of population in England and Wales. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158 (2):203–240, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983290.

Taylor:1995:SAR

[422] Jim Taylor. A statistical analysis of the 1992 research assessment exercise. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158 (2):241–261, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983291.

Korn:1995:ALH

[423] Edward L. Korn and Barry I. Graubard. Analysis of large health surveys: Accounting for the sampling design. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2): 263–295, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983292.

Booth:1995:HDP

[424] Alison L. Booth and Stephen E. Satchell. The hazards of doing a PhD: An analysis of completion and withdrawal rates of British PhD students in the 1980s. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158 (2):297–318, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983293.

Blackburn:1995:MOG

[425] Robert M. Blackburn, Janet Siltanen, and Jennifer Jarman. The measurement of occupational gender segregation: Current problems and a new approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):319–331, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983294.

Dyke:1995:OFY

[426] George Dyke. Obituary: Frank Yates. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158 (2):333–338, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983295.

Plackett:1995:OCE

[427] Robin Plackett. Obituary: Churchill Eisenhart. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):338, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983296.

Renshaw:1995:BRB

[428] Eric Renshaw. Book review: Fractals Everywhere, by M. F. Barnsley. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2): 339–340, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983297.

Lindley:1995:BRB

[429] D. V. Lindley. Book review: Bayesian Theory, by J. M. Bernardo; A. F. M. Smith. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):340–341, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983298.

Davies:1995:BRBc

[430] Neville Davies. Book review: Statistical Methods for SPC and TQM, by D. Bissell. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):341–342, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983299.

Ghani:1995:BRB

[431] Azra Ghani. Book review: AIDS Epidemiology: A Quantitative Approach, by R. Brookmeyer; M. H. Gail. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):343, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983300.

Dewey:1995:BRB

[432] M. Dewey. Book review: Statistical Models in Epidemiology, by D. Clayton; M. Hills. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):343–344, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983301.

Bowman:1995:BRB

[433] Adrian Bowman. Book review: Visualizing Data, by W. S. Cleveland. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2): 344–345, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983302.

Goldstein:1995:BRB

[434] Harvey Goldstein. Book review: Analysis of Longitudinal Data, by P. J. Diggle; K.-Y. Liang; S. L. Zeger. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2): 345–346, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983303.

Marriott:1995:BRB

[435] Paul Marriott. Book review: An Introduction to the Bootstrap, by B. Efron; R. J. Tibshirani. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):347, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983304.

Sheer:1995:BRB

[436] Andrew Sheer. Book review: Competitive Equilibrium: Theory and Applications, by B. Ellickson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2): 347–348, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983305.

Cooper:1995:BRB

[437] R. I. Cooper. Book review: Principles of Applied Statistics, by M. C. Fleming; J. G. Nellis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):348–349, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983306.

Chatfield:1995:BRB

[438] C. Chatfield. Book review: A Handbook of Small Data Sets, by D. J. Hand; F. Daly; A. D. Lunn; K. McConway; E. Ostrowski. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):349, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983307.

Fielding:1995:BRB

[439] Antony Fielding. Book review: Sample Survey Methods and Theory: Vol. 1, Methods and Applications, by M. H. Hansen; W. N. Hurwitz; W. G. Madow; Sample Survey Methods and Theory: Vol. 2, Theory. by M. H. Hansen; W. N. Hurwitz; W. G. Madow. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):350, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983308.

White:1995:BRB

[440] Lynda White. Book review: Design and Analysis of Experiments: Vol. I, Introduction to Experimental Design, by K. Hinkelmann; O. Kempthorne. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):351, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983309.

Lee:1995:BRBa

[441] P. M. Lee. Book review: Simulation and Chaotic Behaviour of α-Stable Stochastic Processes, by A. Janicki; A. Weron. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158 (2):351–352, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983310.

Blower:1995:BRB

[442] G. Blower. Book review: Operator-Limit Distributions in Probability Theory, by Z. J. Jurek; J. D. Mason. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2): 352–353, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983311.

Bingham:1995:BRB

[443] N. H. Bingham. Book review: Probability, Statistics and Optimization: A Tribute to Peter Whittle, by F. P. Kelly. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158 (2):353–354, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983312.

Braithwaite:1995:BRB

[444] G. R. Braithwaite. Book review: Introductory Statistics, by P. S. Mann. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):354, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983313.

Snell:1995:BRBb

[445] Joyce Snell. Book review: Applied Statistics and Probability for Engineers, by D. C. Montgomery; G. C. Runger. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):355, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983314.

Snell:1995:BRBc

[446] Joyce Snell. Book review: Applied Nonparametric Statistical Methods, by P. Sprent. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):355–356, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983315.

Rees:1995:BRB

[447] Mike Rees. Book review: Elementary Probability, by D. Stirzaker. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):356, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983316.

Coleman:1995:BRBa

[448] Rodney Coleman. Book review: Probability Theory, An Analytic View, by D. W. Stroock. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):356–357, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983317.

Walker:1995:BRBb

[449] Monica Walker. Book review: Analysing Offending: Data, Models and Interpretations, by R. Tarling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2): 357–358, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983318.

Anonymous:1995:BMb

[450] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(2):??, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983319.

Anonymous:1995:VI

[451] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3): 639–642, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983436.

Anonymous:1995:FMc

[452] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):??, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983437.

Moore:1995:E

[453] Peter G. Moore. Editorial. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3): 359–361, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983438.

Bartholomew:1995:MUU

[454] David Bartholomew, Peter Moore, Fred Smith, and Paul Allin. The measurement of unemployment in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3): 363–417, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983439.

Chatfield:1995:MUD

[455] Chris Chatfield. Model uncertainty, data mining and statistical inference. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):

419–466, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983440.

McLennan:1995:YCC

[456] Bill McLennan. You can count on us– with confidence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):467–489, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983441.

Nissel:1995:STS

[457] Muriel Nissel. Social trends and social change. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):491–504, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983442.

Day:1995:AID

[458] N. E. Day, S. M. Gore, and D. De Angelis. Acquired Immune Deficiency Syndrome predictions for England and Wales (1992–97): Sensitivity analysis, information, decision. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3): 505–524, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983443.

Bissell:1995:WCS

[459] A. F. Bissell. Weighted cumulative sums for text analysis using word counts. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158 (3):525–545, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983444.

Haberman:1995:EBJ

[460] Steven Haberman and Marlena Schmool.

Estimates of the British Jewish population 1984–88. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):547–562, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983445.

Heath:1995:MCP

[461] Anthony Heath, Geoffrey Evans, and Clive Payne. Modelling the class-party relationship in Britain, 1964–92. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3): 563–574, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983446.

Anonymous:1995:RCS

[462] Anonymous. Report of the Council for the session 1994–95. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3): 575–618, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983447.

Nevill:1995:LE

[463] Alan M. Nevill, Roger L. Holder, Phil McShane, and Richard A. Kronmal. Letters to the editors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):619–625, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983448.

Veevers:1995:ORS

[464] Alan Veevers and Geoff Robinson. Obituary: Richard A. Stone, 1961–94. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3): 627–628, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983449.

Tong:1995:BRB

[465] Howell Tong. Book review: Networks and Chaos–Statistical and Probabilistic Aspects, by O. E. Barndorff-Nielsen; J. L. Jensen; W. S. Kendall. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3): 629–630, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983450.

Atkinson:1995:BRBb

[466] Anthony Atkinson. Book review: Outliers in Statistical Data, by V. Barnett; T. Lewis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):630–631, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983451.

Lievesley:1995:BRB

[467] Denise Lievesley. Book review: Reviews of United Kingdom Statistical Sources: Vol. XXVII, Research and Development, by D. L. Bosworth; R. A. Wilson; A. Young. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):631–632, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983452.

Young:1995:BRB

[468] Karen Young. Book review: The Collected Works of John W. Tukey: Vol. VIII, Multiple Comparisons: 1948– 1983, by John W. Tukey; H. I. Braun. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):632, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983453.

Coutts:1995:BRB

[469] Stewart M. Coutts. Book review: Practical Risk Theory for Actuaries, by C. D. Daykin; T. Pentikainen; M. Pesonen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):633, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983454.

Davies:1995:BRBd

[470] Neville Davies. Book review: A Handbook of Statistical Analysis Using S-Plus, by B. S. Everitt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):633–634, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983455.

Coleman:1995:BRBb

[471] Rodney Coleman. Book review: Studies in Applied Probability: Papers in Honour of Lajos Takacs, by J. Galambos; J. Gani. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):634–635, ???? 1995. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983456.

Ripley:1995:BRBb

[472] B. D. Ripley. Book review: General Pattern Theory: A Mathematical Theory of Regular Structures, by U. Grenander. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):635, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983457.

Davies:1995:BRBe

[473] Neville Davies. Book review: Applied Bayesian Forecasting and Time Series Analysis, by A. Pole; M. West; J. Harrison. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):635–636, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983458.

Whittle:1995:BRB

[474] P. Whittle. Book review: Markov Decision Processes, by M. L. Puterman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158 (3):636–637, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983459.

Lee:1995:BRBb

[475] Peter M. Lee. Book review: Subjective Probability, by G. Wright; P. Ayton. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):637–638, ???? 1995. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983460.

Anonymous:1995:BMc

[476] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 158(3):??, ???? 1995. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983461.

Anonymous:1996:FMa

[477] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):??, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983462.

Evans:1996:EIS

[478] Stephen Evans. Editorial: Issues for statisticians involved with court cases. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):1–3, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983463.

Worcester:1996:PPE

[479] Robert Worcester. Political polling:

95% expertise and 5% luck. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):5–20, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983464.

Lynn:1996:HMO

[480] Peter Lynn and Roger Jowell. How might opinion polls be improved?: The case for probability sampling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1): 21–28, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983465.

Anonymous:1996:DOP

[481] Anonymous. Discussion on opinion polls and general elections. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1): 29–39, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983466.

Inskip:1996:RHV

[482] Hazel M. Inskip. Reay and Hope versus British Nuclear Fuels plc: Issues faced when a research project formed the basis of litigation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):41–47, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983467.

Day:1996:TDB

[483] N. E. Day and S. W. Duffy. Trial design based on surrogate end points — application to comparison of different breast screening frequencies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1): 49–60, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983468.

Ecob:1996:MMA

[484] Russell Ecob. A multilevel modelling approach to examining the effects of area of residence on health and functioning. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):61–75, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983469. See

erratum [672].

Blane:1996:IBM

[485] D. Blane, C. Power, and M. Bartley. Illness behaviour and the measurement of class differentials in morbidity. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):77–92, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983470.

Cook:1996:MCD

[486] Richard J. Cook and Vern T. Farewell.

Multiplicity considerations in the design and analysis of clinical trials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1): 93–110, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983471.

Wadsworth:1996:WME

[487] J. Wadsworth, A. M. Johnson, K. Wellings, and J. Field. What’s in a mean? — an examination of the inconsistency between men and women in reporting sexual partnerships. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):111–123, ????

1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983472.

Lindsey:1996:FBI

[488] J. K. Lindsey. Fitting bivariate intensity functions, with an application to modelling delays in reporting acquired immune deficiency syndrome. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1): 125–131, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983473.

Byron:1996:DTS

[489] R. P. Byron. Diagnostic testing and sensitivity analysis in the construction of social accounting matrices. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1): 133–148, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983474.

Goldstein:1996:UER

[490] Harvey Goldstein and Sally Thomas. Using examination results as indicators of school and college performance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1): 149–163, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983475.

Moore:1996:OJH

[491] Peter G. Moore. Obituary: James Harold Wilson 1916–95. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):165–173, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.jstor.org/
stable/2983476.

Keen:1996:OWJ

[492] Joan Keen. Obituary: William Joseph Jennett, 1912–94. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):173–175, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983477.

Upton:1996:BRB

[493] G. J. G. Upton. Book review: Understanding Log-Linear Analysis with ILOG, by R. Bakeman; B. Robinson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (1):177–178, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983478.

Barnard:1996:BRB

[494] George A. Barnard. Book review:

Inference and Asymptotics, by O. E. Barndorff-Nielsen; D. R. Cox. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1): 178–179, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983479.

Scott:1996:BRB

[495] Marian Scott. Book review: Statistics for the Environment 2: Water Related Issues, by V. Barnett; K. F. Turkman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (1):179–180, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983480.

Ray:1996:BRB

[496] W. D. Ray. Book review: Statistics for Long-Memory Processes, by J. Beran. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):180–181, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983481.

Hidalgo:1996:BRB

[497] J. Hidalgo. Book review: Topics in Advanced Econometrics, by H. J. Bierens. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (1):181–182, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983482.

Snell:1996:BRBa

[498] Joyce Snell. Book review: Alternative Methods of Regression, by D. Birkes; Y. Dodge. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):182–183, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983483.

Jones:1996:BRB

[499] M. C. Jones. Book review: Statistical Models for Ordinal Variables, by C. C. Clogg; E. S. Shihadeh. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1): 183–184, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983484.

Gower:1996:BRB

[500] J. C. Gower. Book review: Multidimensional Scaling, by F. Cox; M. A. A. Cox. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):184–185, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983485.

Jolliffe:1996:BRBa

[501] Flavia Jolliffe. Book review: Teaching Statistics at Its Best: A Series of Articles for Teachers, by D. Green. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1): 185–186, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983486.

Kendall:1996:BRB

[502] Wilfrid Kendall. Book review: Probability with a View Toward Statistics, Vols I and II, by J. Hoffmann-Jorgensen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (1):186–187, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983487.

Constable:1996:BRB

[503] P. D. L. Constable. Book review: Applied Discriminant Analysis, by C. J. Huberty. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):187, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983488.

Snell:1996:BRBb

[504] Joyce Snell. Book review: Modelling Frequency and Count Data, by J. K. Lindsey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):188, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983489.

Dale:1996:BRB

[505] Angela Dale. Book review: Problems and Methods in Longitudinal Research: Stability and Change, by D. Magnusson; L. Bergman; G. Rudinger; B. Torestad. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (1):188–189, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983490.

Woodward:1996:BRB

[506] Mark Woodward. Book review: Analysis of Two-Way Layouts, by J. Mandel. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):189, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983491.

Oman:1996:BRBa

[507] P. Oman. Book review: Analysing Survival Data from Clinical Trials and Observational Studies, by E. Marubini; M. G. Valsecchi. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):190, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983492.

McNicol:1996:BRB

[508] J. W. McNicol. Book review: Agricultural Field Experiments, by R. G. Petersen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):190–191, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983493.

Pitts:1996:BRB

[509] S. M. Pitts. Book review: Theoretical Probability for Applications, by S. C. Port. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):191–192, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983494.

Isham:1996:BRB

[510] Valerie Isham. Book review: Choquet-Deny Type Functional Equations with Applications to Stochastic Models, by C. R. Rao; D. N. Shanbhag. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1): 192–193, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983495.

Walden:1996:BRBa

[511] Andrew Walden. Book review: Statistical Methods for Physical Science: Vol. 28, Methods of Experimental Physics, by J. L. Stanford; S. B. Vardeman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):

193–194, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983496.

Hambly:1996:BRB

[512] Ben Hambly. Book review: Fractals, Random Shapes and Point Fields, by D. Stoyan; H. Stoyan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1): 194–195, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983497.

Anonymous:1996:BMa

[513] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(1):??, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983498.

Anonymous:1996:FMb

[514] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):??, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983166.

Plewis:1996:E

[515] Ian Plewis and Fred Smith. Editorial. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (2):197–199, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983167.

Woodhouse:1996:AME

[516] Geoffrey Woodhouse, Min Yang, Harvey Goldstein, and Jon Rasbash. Adjusting for measurement error in multilevel analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):201–212, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983168.

DeMenezes:1996:NDL

[517] L. M. De Menezes and D. J.

Bartholomew. New developments in latent structure analysis applied to social attitudes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):213–224, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983169.

Pickles:1996:RRD

[518] Andrew Pickles, Kevin Pickering, and Colin Taylor. Reconciling recalled dates of developmental milestones, events and transitions: A mixed generalized linear model with random mean and variance functions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):225–234, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983170.

Simpson:1996:USA

[519] Stephen Simpson, Ian Diamond, Pete Tonkin, and Rachel Tye. Updating small area population estimates in England and Wales. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):235–247, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983171.

Fitzmaurice:1996:LRM

[520] Garrett M. Fitzmaurice, Anthony F. Heath, and Peter Clifford. Logistic regression models for binary panel data with attrition. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):249–263, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983172.

Skinner:1996:RAC

[521] C. J. Skinner and O. Coker. Regression analysis for complex survey data with missing values of a covariate. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2): 265–274, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983173.

Dolton:1996:RER

[522] Peter Dolton and Donal O’Neill. The restart effect and the return to full-time stable employment. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):275–288, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983174.

Steele:1996:IUR

[523] Fiona Steele, Ian Diamond, and Sajeda Amin. Immunization uptake in rural Bangladesh: A multilevel analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):

289–299, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983175.

Francis:1996:VEH

[524] Brian Francis and Mark Fuller. Visualization of event histories. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2): 301–308, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983176.

Smith:1996:MCEa

[525] Peter W. F. Smith, Jonathan J. Forster, and John W. McDonald. Monte Carlo exact tests for square contingency tables. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):309–321, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983177.

Best:1996:BAR

[526] N. G. Best, D. J. Spiegelhalter, A. Thomas, and C. E. G. Brayne. Bayesian analysis of realistically complex models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):323–342, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983178.

Gibson:1996:BRB

[527] Martin Gibson. Book review: Practical Statistical Process Control — A Tool for Quality Manufacturing, by F. Alsup; R. M. Watson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):343–344, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983179.

Young:1996:BRBa

[528] P. J. Young. Book review: The Statistical Analysis of Time Series, by T. W. Anderson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):344–345, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983180.

Kimber:1996:BRB

[529] Alan Kimber. Book review: Practical Methods for Reliability Data Analysis, by J. I. Ansell; M. J. Phillips. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):345, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983181.

Dunn:1996:BRB

[530] Graham Dunn. Book review: Structural Equation Modeling with EQS and EQS/Windows, by B. M. Byrne. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2): 345–346, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983182.

Walden:1996:BRBb

[531] Andrew Walden. Book review: Wavelets: Theory, Algorithms and Applications: Vol. 5, Wavelet Analysis and Its Applications, by C. K. Chui; L. Montefusco; L. Puccio. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):346–347, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983183.

Cooper:1996:BRB

[532] R. I. Cooper. Book review: Elements of Statistics, by F. Daly; D. J. Hand; M. C. Jones; A. D. Lunn; K. J. Mc-Conway. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):347, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983184.

Freeman:1996:BRB

[533] G. H. Freeman. Book review: The Theory of Gambling and Statistical Logic, by R. A. Epstein. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):348, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983185.

Coleman:1996:BRB

[534] Rodney Coleman. Book review: Continuous Univariate Distributions, by N. L. Johnson; S. Kotz; N. Balakrishnan. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):349, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983186.

Jolliffe:1996:BRBb

[535] Flavia Jolliffe. Book review: Just the Essentials of Elementary Statistics, by R. Johnson. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):349–350, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983187.

McKenzie:1996:BRB

[536] E. McKenzie. Book review: Time Series Analysis by Higher Order Crossings, by B. Kedem. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):350, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983188.

White:1996:BRBa

[537] L. V. White. Book review: Quality Through Design: Experimental Design, Off-Line Quality Control, and Taguchi’s Contributions, by N. Logothetis; H. P. Wynn. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):351, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983189.

Knott:1996:BRB

[538] Martin Knott. Book review: Distribution-Free Statistical Methods, by J. S. Maritz. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (2):351–352, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983190.

Tong:1996:BRB

[539] H. Tong. Book review: Chaotic Dynamics: Theory and Applications to Economics, by A. Medio. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2): 352–353, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983191.

Phillips:1996:BRB

[540] Michael J. Phillips. Book review: Limit Theorems of Probability Theory: Sequences of Independent Random Variables, by V. V. Petrov. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2): 353–354, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983192.

Bingham:1996:BRB

[541] N. Bingham. Book review: Diffusions, Markov Processes and Martingales: Vol. 1, Foundations, by L. C. G. Rogers; D. Williams. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):354–355, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983193.

OBroin:1996:BRB

[542] Sean O’Broin. Book review: An Introduction to Statistics for Managers, by A. Spooner; C. Lewis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2): 355–356, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983194.

Lunn:1996:BRB

[543] A. D. Lunn. Book review: Statistical Data Analysis for Ocean and Atmospheric Sciences, by H. J. Thiebaux. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (2):356–357, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983195.

Atkinson:1996:BRBa

[544] A. C. Atkinson. Book review: Advanced Linear Models: Theory and Applications, by S. G. Wang; S.-C. Chow. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (2):357–358, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983196.

Anonymous:1996:BMb

[545] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(2):??, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983197.

Anonymous:1996:VI

[546] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3): 634–638, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983320.

Anonymous:1996:FMc

[547] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):??, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983321.

Lewis:1996:ESS

[548] John A. Lewis. Editorial: Statistics and statisticians in the regulation of medicines. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):359–362, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983322.

Gore:1996:EBS

[549] Sheila M. Gore. Editorial: Bovine spongiform encephalopathy and ‘the term safe is not the same as zero risk’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (3):363–365, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983323.

Smith:1996:MCEb

[550] Adrian F. M. Smith. Mad cows and ecstasy: Chance and choice in an evidencebased society. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):367–383, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983324.

Goldstein:1996:LTT

[551] Harvey Goldstein and David J. Spiegelhalter. League tables and their limitations: Statistical issues in comparisons of institutional performance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3): 385–443, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983325.

Hand:1996:STM

[552] D. J. Hand. Statistics and the theory of measurement. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):445–492, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983326.

Breen:1996:RMU

[553] Richard Breen and Bernadette C. Hayes. Religious mobility in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3): 493–504, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983327.

Goldstein:1996:IAM

[554] Harvey Goldstein and Jon Rasbash. Improved approximations for multilevel models with binary responses. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3): 505–513, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983328.

Congdon:1996:ESL

[555] Peter Congdon. The epidemiology of suicide in London. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):515–533, ????

1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983329.

Smith:1996:POP

[556] T. M. F. Smith. Public opinion polls:

The UK general election, 1992. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3): 535–545, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983330.

Farrington:1996:SAE

[557] C. P. Farrington, N. J. Andrews, A. D.

Beale, and M. A. Catchpole. A statistical algorithm for the early detection of outbreaks of infectious disease. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (3):547–563, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983331.

Anonymous:1996:RCS

[558] Anonymous. Report of the Council for the session 1995–96. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3): 565–610, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983332.

Galbraith:1996:LE

[559] Rex F. Galbraith and Jane I. Galbraith. Letter to the editors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3): 611–613, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983333.

Hendry:1996:OJT

[560] David F. Hendry and Mary S. Morgan. Obituary: Jan Tinbergen, 1903– 94. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (3):614–616, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983334.

Bather:1996:OSV

[561] John Bather. Obituary: Stefan Vajda, 1901–95. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):616–617, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983335.

Muba:1996:OPP

[562] V. E. Muba and G. M. Clarke. Obituary: Pius Peter Kallaa, 1944–95. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3): 617–618, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983336.

Disney:1996:BRB

[563] John Disney. Book review: Quality by Experimental Design, by T. B. Barker. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (3):619–620, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983337.

Kemp:1996:BRB

[564] David Kemp. Book review: Statistics for the Environment, by V. Barnett; K. F. Turkman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):620–621, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983338.

Forster:1996:BRB

[565] Jon Forster. Book review: Aspects of Uncertainty: A Tribute to D. V. Lindley, by D. V. Lindley; P. R. Freeman; A. F. M. Smith. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):621, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983339.

Fielding:1996:BRB

[566] Antony Fielding. Book review: Experimental Methods: A Primer for Economists, by D. Friedman; S. Sunder. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):622, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983340.

Rennolls:1996:BRB

[567] Keith Rennolls. Book review: Statistical Inference, by P. H. Garthwaite; I. T. Jolliffe; B. Jones. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):622–623, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983341.

Atkinson:1996:BRBb

[568] Rachel Atkinson. Book review: An Introduction to Experimental Design and Statistics for Biology, by D. Heath. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (3):623–624, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983342.

Lee:1996:BRB

[569] P. M. Lee. Book review: Random Walks and Random Environments: Vol. I, Random Walks, by B. D. Hughes. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (3):624–625, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983343.

White:1996:BRBb

[570] Lynda White. Book review: Cyclic and Computer Generated Designs, by J. A. John; E. R. Williams. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):625, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983344.

Ringrose:1996:BRB

[571] Trevor J. Ringrose. Book review: Recent Advances in Descriptive Multivariate Analysis, by W. J. Krzanowski. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (3):625–626, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983345.

Winterbottom:1996:BRB

[572] Alan Winterbottom. Book review: Reliability Analysis for Engineers: An Introduction, by R. D. Leitch. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3): 626–627, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983346.

Titterington:1996:BRB

[573] D. M. Titterington. Book review:

Stochastic Geometry Models in Image Analysis and Spatial Statistics, by M. N. M. van Lieshout. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):627–628, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983347.

Hinde:1996:BRB

[574] John Hinde. Book review: Statistical Applications Using Fuzzy Sets, by K. G. Manton; M. A. Woodbury; H. D. Tolley. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (3):628–629, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983348.

Braithwaite:1996:BRB

[575] G. R. Braithwaite. Book review: Statistics in Engineering — A Practical Approach, by A. V. Metcalfe. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3): 629–630, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983349.

Oman:1996:BRBb

[576] P. Oman. Book review: Survival Analysis — A Practical Approach, by M. K. B. Parmar; D. Machin. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3): 630–631, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983350.

Young:1996:BRBb

[577] G. A. Young. Book review: The Jackknife and Bootstrap, by J. Shao; D. Tu. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159 (3):631–632, ???? 1996. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983351.

Orchard:1996:BRB

[578] T. J. Orchard. Book review: Business Forecasting and Planning, by P. Shearer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):632, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983352.

Lound:1996:BRB

[579] Charles Lound. Book review: Statistics for the Social Sciences, by R. M. Sirkin. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):632–633, ???? 1996. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983353.

Anonymous:1996:BMc

[580] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 159(3):??, ???? 1996. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983354.

Anonymous:1997:FMa

[581] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):??, ???? 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983376.

Roberts:1997:EHS

[582] Dennis Roberts. Editorial: Harmonization of statistical definitions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):1–4, ???? 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983377.

Steel:1997:PME

[583] David Steel. Producing monthly estimates of unemployment and employment according to the International Labour Office definition. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):5–46, ???? 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983378.

Wright:1997:CSM

[584] Eileen M. Wright and Patrick Royston. A comparison of statistical methods for age-related reference intervals.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (1):47–69, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983379.

Alho:1997:SUC

[585] Juha M. Alho. Scenarios, uncertainty and conditional forecasts of the world population. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):71–85, ???? 1997. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983380.

Anderson:1997:SMI

[586] N. H. Anderson and D. M. Titterington. Some methods for investigating spatial clustering, with epidemiological applications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):87–105, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983381.

Davies:1997:ICD

[587] Hugh Davies, Heather Joshi, and Lynda Clarke. Is it cash that the deprived are short of? Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):107–126, ???? 1997. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983382.

Brookmeyer:1997:AFB

[588] Ron Brookmeyer. Accounting for followup bias in estimation of human immunodeficiency virus incidence rates. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1): 127–140, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983383.

Watts:1997:LEM

[589] Martin Watts. Letter to the editor:

The measurement of occupational gender segregation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):141–145, ???? 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Watts:1997:LE

[590] Martin Watts, R. M. Blackburn, J. Jarman, and J. Siltanen. Letter to the editors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):141–147, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983384.

Rudoe:1997:OSH

[591] W. Rudoe. Obituary: Sir Harry Campion, 1905–96. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):148–151, ???? 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983385.

Smith:1997:OMC

[592] Cedric A. B. Smith. Obituary: Maurice Charles Kenneth Tweedie, 1919– 96. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (1):151–154, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983386.

Lambert:1997:BRB

[593] J. A. Lambert. Book review: Statistics and the Evaluation of Evidence for Forensic Scientists, by C. G. G. Aitken. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (1):155–156, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983387.

Anonymous:1997:BR

[594] Anonymous. Book reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1): 155–164, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Castell:1997:BRB

[595] Paul Castell. Book review: Uncertain Belief: Is it Rational to be a Christian?, by D. J. Bartholomew. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):156, ???? 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983388.

Bather:1997:BRB

[596] John Bather. Book review: A Mathematical Kaleidoscope: Applications in Industry, Business and Science, by B. Conolly; S. Vajda. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):157, ???? 1997. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983389.

Hionidou:1997:BRB

[597] Violetta Hionidou. Book review: A New Social Atlas of Britain, by D. Dorling. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (1):157–158, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983390.

Winstanley:1997:BRB

[598] Bridget Winstanley. Book review: International Statistics Sources: Subject Guide to Sources of International Comparative Statistics. Vol 1, 2, by M. C. Fleming; J. G. Nellis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):158, ???? 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983391.

Stander:1997:BRB

[599] Julian Stander. Book review: Markov Chain Monte Carlo in Practice, by W. R. Gilks; S. Richardson; D. J. Speigelhalter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):158–159, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983392.

Lee:1997:BRB

[600] Peter M. Lee. Book review: Monotone Structure in Discrete Event Systems, by P. Glasserman; D. D. Yao. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (1):159–160, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983393.

Hartley:1997:BRB

[601] R. Hartley. Book review: Discrete-time Controlled Markov Processes: Basic Optimality Criteria, by O. Hernández-Lerna; J. B. Lasserre. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):160–161, ???? 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2983394.

Moustaki:1997:BRB

[602] Irini Moustaki. Book review: Advanced Structural Equation Modelling: Issues and Techniques, by G. A. Marcoulides; R. E. Schumacker. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1): 161–162, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983395.

Braithwaite:1997:BRB

[603] G. R. Braithwaite. Book review: Linear Models: Least Squares and Alternatives, by C. R. Rao; H. Toutenberg. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):162, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983396.

Rouncefield:1997:BRB

[604] Mary Rouncefield. Book review: Advanced Modular Mathematics: Pure Mathematics 1, by G. Smithers; Advanced Modular Mathematics: Statistics 1 by G. Westover. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1): 163–164, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983397.

Kent:1997:BRB

[605] John T. Kent. Book review: The Geometry of Multivariate Statistics, by T. D. Wickens. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):164, ???? 1997. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983398.

Anonymous:1997:BMa

[606] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(1):??, ???? 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983399.

Nanson:1997:ETN

[607] Bryan Nanson. Editorial: Towards a national geospatial data framework. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2): 165–167, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983212.

Moore:1997:DUN

[608] Peter G. Moore. The development of the UK national lottery: 1992–96. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2): 169–185, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983213. See [611].

Haigh:1997:SNL

[609] John Haigh. The statistics of the National Lottery. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):187–206, ???? 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/

2983214. See [611].

King:1997:PNL

[610] J. R. B. King. Participation in the National Lottery — evidence from the Family Expenditure Survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2): 207–212, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983215.

Anonymous:1997:DPM

[611] Anonymous. Discussion of the papers by Moore and Haigh. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2): 213–235, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983216. See [608,

609].

Copas:1997:ERR

[612] J. B. Copas and F. Heydari. Estimating the risk of reoffending by using exponential mixture models. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):

237–252, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983217.

Lynn:1997:SFE

[613] Peter Lynn. Sampling frame effects on the British Crime Survey. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2): 253–269, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983218.

Aitken:1997:TIC

[614] C. G. G. Aitken. The third International Conference on Forensic Statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (2):271–273, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983219.

Finkelstein:1997:CCG

[615] Michael O. Finkelstein and Bruce Levin.

Clear choices and guesswork in peremptory challenges in Federal criminal trials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (2):275–288, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983220.

Gastwirth:1997:SED

[616] Joseph L. Gastwirth. Statistical evidence in discrimination cases. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2): 289–303, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983221.

Mellen:1997:MSD

[617] Beverly G. Mellen and Richard M. Royall. Measuring the strength of deoxyribonucleic acid evidence, and probabilities of strong implicating evidence. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2): 305–320, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983222.

Fienberg:1997:EEW

[618] Stephen E. Fienberg. Ethics and the expert witness: Statistics on trial. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2): 321–331, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983223.

Aitken:1997:EQD

[619] C. G. G. Aitken, J. Bring, T. Leonard, and O. Papasouliotis. Estimation of quantities of drugs handled and the burden of proof. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):333–350, ???? 1997. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983224.

Kadane:1997:MDF

[620] Joseph B. Kadane and Norma Terrin. Missing data in the forensic context. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (2):351–357, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983225.

Knolle:1997:LE

[621] Helmut Knolle, Jane Wadsworth, and Anne M. Johnson. Letter to the editors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):359–361, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983226.

Suhov:1997:OBG

[622] Yurii Suhov. Obituary: Boris Gnedenko, 1912–95. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):363–366, ???? 1997. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983227.

Winterbottom:1997:BRB

[623] Alan Winterbottom. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):367–368, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983228.

Anonymous:1997:BRRa

[624] Anonymous. Book review: Reviews.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (2):367–380, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Bowie:1997:BRB

[625] D. C. Bowie. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):369, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983229.

Bingham:1997:BRBa

[626] N. H. Bingham. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):370, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983230.

Hitchcock:1997:BRB

[627] S. E. Hitchcock. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):370–371, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983231.

Orriss:1997:BRB

[628] John Orriss. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (2):371–372, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983232.

Marriott:1997:BRBa

[629] Paul Marriott. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):372, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983233.

Joshi:1997:BRB

[630] Heather Joshi. Book review: [untitled].

Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (2):372–373, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983234.

Kendall:1997:BRBa

[631] Wilfrid Kendall. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):373–374, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983235.

Coleman:1997:BRB

[632] Rodney Coleman. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):374–375, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983236.

Collett:1997:BRB

[633] Dave Collett. Book review: [untitled].

Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (2):375–376, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983237.

Bingham:1997:BRBb

[634] N. H. Bingham. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):376–377, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983238.

Evett:1997:BRB

[635] Ian W. Evett. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (2):377–378, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983239.

Marriott:1997:BRBb

[636] Paul Marriott. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (2):378–379, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983240.

Marriott:1997:BRBc

[637] Paul Marriott. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):379, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983241.

Hidalgo:1997:BRB

[638] Javier Hidalgo. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):380, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983242.

Anonymous:1997:BMb

[639] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):??, ????

1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983243.

Anonymous:1997:FMb

[640] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(2):??, ???? 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983211.

OMuircheartaigh:1997:EUP

[641] Colm O’Muircheartaigh and Peter Lynn.

Editorial: The 1997 UK pre-election polls. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):381–385, September 1997. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983261.

Bartholomew:1997:EMU

[642] David Bartholomew. Editorial: The measurement of unemployment in the UK: The position at June 1997. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3): 385–388, September 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983262.

Chesher:1997:DRS

[643] Andrew Chesher. Diet revealed?:

Semiparametric estimation of nutrient intake–age relationships. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):389– 428, September 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983263.

Foreman:1997:BAD

[644] Lindsey A. Foreman, Adrian F. M. Smith, and Ian W. Evett. Bayesian analysis of DNA profiling data in forensic identification applications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3): 429–469, September 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983264.

Byers:1997:MPP

[645] David Byers, James Davidson, and David Peel. Modelling political popularity: An analysis of long-range dependence in opinion poll series. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3): 471–490, September 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983265. See cor-

rection [760].

Diggle:1997:RMD

[646] Peter Diggle, Sara Morris, Paul Elliott, and Gavin Shaddick. Regression modelling of disease risk in relation to point sources. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):491–505, September 1997. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2983266.

Carpenter:1997:EAB

[647] Lucy M. Carpenter, Noreen E. S. Maconochie, Eve Roman, and D. R. Cox.

Examining associations between occupation and health by using routinely collected data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):507–521, September 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983267.

Hand:1997:SCM

[648] D. J. Hand and W. E. Henley. Statistical classification methods in consumer credit scoring: A review. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3): 523–541, September 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983268.

Anonymous:1997:MRR

[649] Anonymous. Meeting report: Report of the Council for the session 1996– 97, presented at the one hundred and sixty-third Annual General Meeting of the Royal Statistical Society, June 18th, 1997. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):543–592, September 1997. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983269.

Longford:1997:LE

[650] N. T. Longford. Letter to the editors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):593, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983270.

Longford:1997:LEI

[651] N. T. Longford. Letter to the editors: Improved approximations for multilevel models with binary responses. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3): 593, September 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Thatcher:1997:ORF

[652] Roger Thatcher. Obituary: Ronald Frederick Fowler, 1910–97. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):594– 595, September 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983271.

Oman:1997:BRB

[653] P. Oman. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3): 596–597, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983272.

Anonymous:1997:BRRb

[654] Anonymous. Book review: Reviews. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3): 596–608, September 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Collop:1997:BRB

[655] M. J. Collop. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):597, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983273.

Sweeting:1997:BRB

[656] Trevor J. Sweeting. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):598, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983274.

Baxter:1997:BRB

[657] Martin Baxter. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):598–599, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983275.

Grubb:1997:BRB

[658] Howard Grubb. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):599–600, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983276.

Pullinger:1997:BRB

[659] John Pullinger. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):600–601, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983277.

Donev:1997:BRB

[660] A. N. Donev. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (3):601–602, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983278.

Jolliffe:1997:BRB

[661] Flavia Jolliffe. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (3):602–603, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983279.

Titterington:1997:BRB

[662] D. M. Titterington. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):603–604, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983280.

McCrudden:1997:BRB

[663] Mick McCrudden. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):604, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983281.

Fielding:1997:BRB

[664] Antony Fielding. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):604–605, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983282.

Cox:1997:BRB

[665] D. R. Cox. Book review: [untitled].

Journal of the Royal Statistical Society. Series A (Statistics in Society), 160 (3):605–606, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983283.

Thomson:1997:BRB

[666] Norman Thomson. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):606–607, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983284.

Bingham:1997:BRBc

[667] Michael S. Bingham. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):607, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983285.

Kendall:1997:BRBb

[668] David Kendall. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):607–608, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983286.

Marriott:1997:BRBd

[669] Paul Marriott. Book review: [untitled]. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):608, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983287.

Anonymous:1997:IAI

[670] Anonymous. Index: Author index: Volume 160, 1997. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):609–612, September 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:1997:VI

[671] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3): 609–612, ???? 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983259.

Ecob:1997:ECM

[672] Russell Ecob. Erratum: Corrigendum: A multilevel modelling approach to examining the effects of area of residence on health and functioning. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3): 612, September 1997. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [484].

Anonymous:1997:BMc

[673] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):??, ???? 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983288.

Anonymous:1997:FMc

[674] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 160(3):??, ???? 1997. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983260.

Plewis:1998:ELF

[675] Ian Plewis and Gillian Raab. Editorial:

Looking forwards. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(1):1–2, ???? 1998. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983549.

Rose:1998:EES

[676] David Rose and Karen O’Reilly. Editorial: The Economic and Social Research Council review of government social classifications. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(1):2–4, ???? 1998. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983550.

Deely:1998:QRC

[677] John J. Deely and Adrian F. M. Smith.

Quantitative refinements for comparisons of institutional performance. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161 (1):5–12, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983551.

Pfeffermann:1998:EGF

[678] Danny Pfeffermann, Chris Skinner, and Keith Humphreys. The estimation of gross flows in the presence of measurement error using auxiliary variables. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161 (1):13–32, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983552.

Davies:1998:GII

[679] Hugh Davies and Heather Joshi. Gender and income inequality in the UK 1968– 1990: The feminization of earnings or of poverty? Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(1):33–61, ???? 1998. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983553.

OMuircheartaigh:1998:RII

[680] Colm O’Muircheartaigh and Pamela Campanelli. The relative impact of interviewer effects and sample design effects on survey precision. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(1): 63–77, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983554.

Royston:1998:MEA

[681] Patrick Royston and Eileen M. Wright. A method for estimating age-specific reference intervals (‘normal ranges’) based on fractional polynomials and exponential transformation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(1):79–101, ???? 1998. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983555.

Fatti:1998:BUR

[682] L. P. Fatti, E. M. Senaoana, and M. L. Thompson. Bayesian updating in reference centile charts. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(1): 103–115, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983556.

Anonymous:1998:BMa

[683] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(1):??, ???? 1998. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983557.

Anonymous:1998:FMa

[684] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(1):??, ???? 1998. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983548.

Raab:1998:EDB

[685] Gillian Raab and Uma Moorthy. Editorial: Double-blind refereeing. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(2): 117–119, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983245.

Langford:1998:OMD

[686] Ian H. Langford and Toby Lewis. Outliers in multilevel data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(2): 121–160, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983246.

Anderson:1998:RSH

[687] Roy M. Anderson and Sarah C. Darby. The role of statistics in human immunodeficiency virus research. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(2): 161–166, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983247.

DeAngelis:1998:AID

[688] D. De Angelis, N. E. Day, and O. N. Gill. Acquired immune deficiency syndrome projections in England and Wales: Interplay of methodology and data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(2): 167–176, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983248.

Sabin:1998:CMD

[689] Caroline A. Sabin, Amanda Mocroft, Alessandro Cozzi Lepri, and Andrew N. Phillips. Cofactors and markers of disease progression in human immunodeficiency virus infection. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(2): 177–189, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983249.

Darby:1998:MMS

[690] Sarah C. Darby and David W. Ewart.

Measures of mortality in the study of individuals infected with the human immunodeficiency virus in the UK haemophilia population. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(2): 191–198, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983250.

Holmes:1998:HIV

[691] E. C. Holmes. Human immunodeficiency virus, DNA and statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(2): 199–208, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983251.

Garnett:1998:IBH

[692] Geoff P. Garnett. The influence of behavioural heterogeneity on the population level effect of potential prophylactic type 1 human immunodeficiency virus vaccines. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(2):209–225, ???? 1998. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983252.

Ghani:1998:MSP

[693] A. C. Ghani and G. P. Garnett. Measuring sexual partner networks for transmission of sexually transmitted diseases.

Journal of the Royal Statistical Society. Series A (Statistics in Society), 161 (2):227–238, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983253.

Babiker:1998:IDA

[694] Abdel G. Babiker, Janet H. Darbyshire, Tim E. A. Peto, and A. Sarah Walker. Issues in the design and analysis of therapeutic trials in human immunodeficiency virus infection. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(2): 239–249, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983254.

Hayes:1998:DHI

[695] Richard Hayes. Design of human immunodeficiency virus intervention trials in developing countries. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(2): 251–263, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983255.

Halloran:1998:SIP

[696] M. Elizabeth Halloran. Statistical issues in phase III human immunodeficiency virus vaccine trial design. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(2): 265–272, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983256.

Kulldorff:1998:LE

[697] Martin Kulldorff. Letter to the editors. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(2):273, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983257.

Anonymous:1998:BMb

[698] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(2):??, ???? 1998. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983258.

Anonymous:1998:FMb

[699] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(2):??, ???? 1998. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983244.

Curnow:1998:EIN

[700] Robert Curnow. Editorial: An independent national statistical service. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(3): 275–277, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983200.

Allin:1998:ELM

[701] Paul Allin. Editorial: Labour market statistics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(3):277–280, ???? 1998. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983201.

Custance:1998:SID

[702] John Custance and Hilary Hillier. Statistical issues in developing indicators of sustainable development. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(3): 281–290, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983202.

Levett:1998:SII

[703] Roger Levett. Sustainability indicators — integrating quality of life and environmental protection. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(3):291–302, ???? 1998. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983203.

Barnett:1998:DMA

[704] Vic Barnett. Discussion at the meeting on ‘Alternatives to Economic Statistics as Indicators of National Well-Being’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161 (3):303–311, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983204.

Wakeford:1998:ELS

[705] Richard Wakeford. Epidemiology and litigation — the Sellafield childhood leukaemia cases. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(3):313–325, ???? 1998. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/

2983205. See corrigendum [723].

Bell:1998:WTG

[706] David N. F. Bell and Robert A. Hart.

Working time in Great Britain, 1975– 1994: Evidence from the New Earnings Survey Panel data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(3):327–348, ???? 1998. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983206.

Mohamed:1998:DIM

[707] Wan Norsiah Mohamed, Ian Diamond, and Peter W. F. Smith. The determinants of infant mortality in Malaysia: A graphical chain modelling approach. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161 (3):349–366, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2983207.

Raab:1998:MHI

[708] G. M. Raab, S. Yang, G. Allardice, D. J.

Goldberg, and J. McMenamin. Modelling human immunodeficiency virus infection and acquired immune deficiency syndrome cases in Scotland: Data sources, prior information and Bayesian estimation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(3):367–384, ???? 1998. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2983208.

Copas:1998:DNI

[709] Andrew J. Copas and Vern T.

Farewell. Dealing with non-ignorable non-response by using an ‘enthusiasmto-respond’ variable. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(3):385–396, ???? 1998. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2983209.

Anonymous:1998:AII

[710] Anonymous. Author index: Index of authors, volume 161, 1998. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(3): 397–398, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:1998:VI

[711] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(3): 397–398, ???? 1998. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2983198.

Anonymous:1998:CCV

[712] Anonymous. Contents: Contents of volume 161, 1998. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(3):399–400, ???? 1998. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:1998:BMc

[713] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(3):??, ???? 1998. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2983210.

Anonymous:1998:FMc

[714] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 161(3):??, ???? 1998. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2983199.

Collins:1999:ESU

[715] Martin Collins. Editorial: Sampling for UK telephone surveys. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(1):1–4, ???? 1999. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2680464.

Thatcher:1999:LTP

[716] A. R. Thatcher. The long-term pattern of adult mortality and the highest attained age. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(1):5–43, ???? 1999. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680465.

Spiegelhalter:1999:SAS

[717] David J. Spiegelhalter. Surgical audit:

Statistical lessons from Nightingale and Codman. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(1):45–58, ???? 1999. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680466.

Blane:1999:DSM

[718] D. Blane, S. Harding, and M. Rosato.

Does social mobility affect the size of the socioeconomic mortality differential?: Evidence from the office for national statistics longitudinal study. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162 (1):59–70, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680467.

Sauerbrei:1999:BMP

[719] W. Sauerbrei and P. Royston. Building multivariable prognostic and diagnostic models: Transformation of the predictors by using fractional polynomials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(1):71–94, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680468. See corri-

gendum [761].

Copas:1999:WWS

[720] John Copas. What works?: Selectivity models and meta-analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(1): 95–109, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680469.

Firth:1999:EPU

[721] David Firth, Clive Payne, and Joan Payne. Efficacy of programmes for the unemployed: Discrete time modelling of duration data from a matched comparison study. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(1):111–120, ???? 1999. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680470.

ONeill:1999:BIP

[722] Philip D. O’Neill and Gareth O.

Roberts. Bayesian inference for partially observed stochastic epidemics. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(1): 121–129, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680471.

Wakeford:1999:CEL

[723] R. Wakeford. Corrigendum: Epidemiology and litigation — the Sellafield childhood leukaemia cases. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(1):131, ???? 1999. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/

2680472. See [705].

Anonymous:1999:BMa

[724] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(1):??, ???? 1999. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2680473.

Anonymous:1999:FMa

[725] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(1):??, ???? 1999. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2680463.

McLean:1999:ESD

[726] Iain McLean. Editorial: Statistics and devolution in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(2): 133–136, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680575.

Broughton:1999:ILL

[727] J. Broughton, M. Hazelton, and M. Stone. Influence of light level on the incidence of road casualties and the predicted effect of changing ‘summertime’. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(2):137–175, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680576.

OMuircheartaigh:1999:SPM

[728] Colm O’Muircheartaigh and Irini Moustaki. Symmetric pattern models: A latent variable approach to item nonresponse in attitude scales. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(2): 177–194, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680577.

Bohning:1999:ZIP

[729] Dankmar Böhning, Ekkehart Dietz, Peter Schlattmann, Lisette Mendonça, and Ursula Kirchner. The zero-inflated Poisson model and the decayed, missing and filled teeth index in dental epidemiology. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(2):195–209, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680578. See corri-

gendum [759].

Brown:1999:FBE

[730] P. J. Brown, D. Firth, and C. D. Payne.

Forecasting on British Election Night 1997. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(2):211–226, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680579.

Longford:1999:MSE

[731] Nicholas T. Longford. Multivariate shrinkage estimation of small area means and proportions. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(2):227–245, ???? 1999. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/

2680580. See discussion [740].

Brown:1999:MSO

[732] J. J. Brown, I. D. Diamond, R. L.

Chambers, L. J. Buckner, and A. D. Teague. A methodological strategy for a one-number census in the UK. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(2): 247–267, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680581.

Anonymous:1999:BMb

[733] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(2):??, ????

1999. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2680582.

Anonymous:1999:FMb

[734] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(2):??, ???? 1999. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/
                                        2680574.

Plewis:1999:EMS

[735] Ian Plewis and Gillian Raab. Editorial: Modelling structured categorical data. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):269–271, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680476.

Diamond:1999:SVT

[736] Ian Diamond, Steve Clements, Nicole Stone, and Roger Ingham. Spatial variation in teenage conceptions in South and West England. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):273–289, ???? 1999. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/stable/

2680477. See discussion [740].

Longford:1999:EPH

[737] Nicholas T. Longford. An experiment in primary health care provision. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3): 291–302, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680478.

Fielding:1999:WUA

[738] Antony Fielding. Why use arbitrary points scores?: Ordered categories in models of educational progress. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3): 303–328, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680479. See dis-

cussion [740].

Payne:1999:CPD

[739] C. D. Payne. Comments on the papers by Diamond et al., Longford and Fielding. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):329–330, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). See [736, 737, 738].

Steele:1999:CPD

[740] Fiona Steele. Comments on the papers by Diamond et al., Longford and Fielding. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):329–330, ???? 1999. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680480. See

[736, 737, 738].

Crouchley:1999:CPA

[741] R. Crouchley and R. B. Davies. A comparison of population average and random-effect models for the analysis of longitudinal count data with base-line information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):331–347, ???? 1999. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2680481. See

comments [743].

Ribaudo:1999:MAL

[742] Heather J. Ribaudo, Marisa Bacchi, Jurg Bernhard, and Simon G. Thompson. A multilevel analysis of longitudinal ordinal data: Evaluation of the level of physical performance of women receiving adjuvant therapy for breast cancer. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162 (3):349–360, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680482. See [743].

Goldstein:1999:CPC

[743] Harvey Goldstein. Comments on the papers by Crouchley and Davies and Ribaudo et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):361, ???? 1999. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2680483. See

[741, 742].

Payne:1999:CPCa

[744] C. D. Payne. Comments on the papers by Crouchley and Davies and Ribaudo et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):361, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Best:1999:AIP

[745] Nicky Best and Jon Wakefield. Accounting for inaccuracies in population counts and case registration in cancer mapping studies. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):363–382, ???? 1999. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680484. See

discussion [748].

Fienberg:1999:CMB

[746] Stephen E. Fienberg, Matthew S. Johnson, and Brian W. Junker. Classical multilevel and Bayesian approaches to population size estimation using multiple lists. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):383–405, ???? 1999. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680485. See

discussion [748].

Romaniuk:1999:MCU

[747] H. Romaniuk, C. J. Skinner, and P. J.

Cooper. Modelling consumers’ use of products. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):407–421, ???? 1999. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

www.jstor.org/stable/2680486. See

discussion [748].

Jones:1999:CPB

[748] David R. Jones. Comments on the papers by Best and Wakefield, Fienberg et al. and Romaniuk et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):423, ???? 1999. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/

2680487. See [745, 746, 747].

Payne:1999:CPB

[749] C. D. Payne. Comments on the papers by Best and Wakefield, Flenberg et al. and Romaniuk et al. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):423, ???? 1999. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Clayton:1999:ELC

[750] David Clayton and Jon Rasbash. Estimation in large crossed random-effect models by data augmentation. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3): 425–436, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/2680488.

OMuircheartaigh:1999:MER

[751] Colm O’Muircheartaigh and Pamela Campanelli. A multilevel exploration of the role of interviewers in survey nonresponse. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):437–446, ???? 1999. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        www.jstor.org/stable/2680489. See

[750].

Payne:1999:CPCb

[752] C. D. Payne. Comments on the papers by Clayton and Rasbash, and O’Muircheartaigh and Campanelli. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162 (3):447–448, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Smith:1999:CPC

[753] Peter W. F. Smith. Comments on the papers by Clayton and Rasbash, and O’Muircheartaigh and Campanelli. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162 (3):447–448, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680490. See [750,

751].

Anonymous:1999:IIA

[754] Anonymous. Index: Index of authors, volume 162, 1999. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):449–450, ???? 1999. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:1999:VI

[755] Anonymous. Volume information. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3): 449–452, ???? 1999. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/2680474.

Anonymous:1999:ICV

[756] Anonymous. Index: Contents of volume 162, 1999. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):450–451, ???? 1999. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

Anonymous:1999:BMc

[757] Anonymous. Back matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):??, ????

1999. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2680491.

Anonymous:1999:FMc

[758] Anonymous. Front matter. Journal of the Royal Statistical Society. Series A (Statistics in Society), 162(3):??, ???? 1999. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/
2680475.

Anonymous:2000:CZI

[759] Anonymous. Corrigendum: The zeroinflated Poisson model and the decayed, missing and filled teeth index in dental epidemiology. Journal of the Royal Statistical Society. Series A (Statistics in Society), 163(1):121–122, ???? 2000. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/

2680512. See [729].

Byers:2002:MPP

[760] David Byers, James Davidson, and David Peel. Modelling political popularity: A correction. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(1):187–189, ???? 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/

3559768. See [645].

Sauerbrei:2002:CBM

[761] W. Sauerbrei and P. Royston. Corrigendum: Building multivariable prognostic and diagnostic models: Transformation of the predictors by using fractional polynomials. Journal of the Royal Statistical Society. Series A (Statistics in Society), 165(2):399–400, June 2002. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/stable/

3559934. See [719].